### v7

# Spatial Scale of Agglomeration and Dispersion Number, Spacing, and the Spatial Extent of Cities

Spatial Scale of Agglomeration and Dispersion
Number, Spacing, and the Spatial Extent of Cities
Takashi Akamatsu Tomoya Mori Tohoku Univ. Kyoto Univ., RIETI Minoru Osawa Yuki Takayama Kyoto Univ. Science Tokyo TA: akamatsu@plan.civil.tohoku.ac.jp, TM: mori@kier.kyoto-u.ac.jp, MO: osawa.minoru.4z@kyoto-u.ac.jp, YT: takayama.y.cc65@m.isct.ac.jp. We are indebted to Treb Allen, Kristian Behrens, Donald Davis, Gilles Duranton, Yasusada Murata, Jonathan Newton, Michael Pflüger, Diego Puga, Esteban Rossi-Hansberg, Shunsuke Segi, Akihisa Shibata, and Jacques-François Thisse for their constructive comments. We are grateful to anonymous referees and to all seminar and conference participants. This research has been supported by JSPS KAKENHI 17H00987, 18H01556, 19K15108, 21K04299, 22K04353, 23K22880, 23K22887, 24K00999, and 25H00543. This research was conducted as part of the project, “Agglomeration-based Framework for Empirical and Policy Analyses of Regional Economies,” undertaken at the Research Institute of Economy, Trade and Industry. We acknowledge financial support from the Kajima Foundations and the Murata Science Foundation.
(July 4, 2025)
Cities around the world have become fewer, larger, and more spatially dispersed. We offer a theoretical framework that can account for this dual transformation—economy-wide concentration and local spreading—by distinguishing between two types of dispersion forces: “local” forces acting within cities, and “global” forces acting across them. The distinction leads to a systematic classification of spatial models into a few fundamental types, each entailing distinct spatial patterns and comparative statics. The framework serves to reconcile divergent empirical findings and clarifies how spatial outcomes of transport policies may depend on the spatial scale at which relevant dispersion forces operate.
JEL: C62, R12, R13
Keywords: agglomeration patterns; spatial scale of dispersion forces, many regions, economic geography
1   Introduction
Urban development often exhibits patterns that appear contradictory at first glance. In Japan, the post-1960 construction of nationwide high-speed railway and highway networks [^1] Between 1970 and 2020, the total length increased from 1,119 km to 9,050 km for highways, and from 515 km to 3,106 km for high-speed railways—an increase of over eightfold and sixfold, respectively. See Fig. B.1 in Appendix B. took place alongside rapid demographic shifts. Over the subsequent five decades, urban populations became increasingly concentrated in larger cities: the population share of Japan’s top 100 cities rose by 19 %, while that of the remaining cities fell by 17 %, as reflected in the steeper slope of the rank–size plot (Fig. 1A). At the same time, however, population became more spatially dispersed within each city: between 1970 and 2020, the average city experienced a 35 % decline in its maximum population density and a 24 % decline in average density (Fig. 1B). Local spreading is also evident in within-city population distributions (Fig. 1C). The co-occurrence of economy-wide concentration and local spreading has also been observed in countries such as China, France, and the United States over the recent decades. [^2] See Appendix B. Combes-et-al-DP2023 obtained similar results for France for a much longer duration of time by constructing original panel data for French cities during the 1760–2020 period. The recurring dual pattern across varied institutional and historical settings points to a deeper mechanism operating across spatial scales.
A Population rank–size plot
B Maximum and average population density
C The population distribution within Tokyo in 1970 (left) and 2020 (right)
Figure 1: Global concentration and local dispersion in Japan from 1970 to 2020.
Note: A city is defined as a cluster of contiguous 1km-by-1km grid cells, each with a population density of at least 1,000/km2 and collectively comprising at least 10,000 residents (see Appendix B for details). The number of identified cities decreased from 504 in 1970 to 431 in 2020. Panel (B) shows the annual cross-city arithmetic means of maximum and average population densities, along with the 90th percentile band. Panel (C) visualizes the within-city population distribution in Tokyo; similar figures for Osaka (Japan’s second-largest city) and Nagoya (third-largest) are provided in Figs. 4A and 4B.
D Full agglomeration
E Monocentric distribution
F Polycentric distribution
Figure 2: Spatial distributions in a square economy with uniform local fundamentals.
Motivated by multi-scale agglomeration and dispersion phenomena, this study develops a conceptual framework to distinguish between “local” and “global” dispersion forces, each acting at a distinct spatial scale. While many spatial models implicitly embed local or global dispersion forces, we formalize the distinction that has received limited systematic attention in the literature. Our framework follows the literature on spatial agglomeration to employ stylized geographical settings such as a symmetric two-region economy. This allows us to provide systematic comparison of endogenous mechanisms across a wide range of spatial models, including canonical quantitative models such as Redding-Sturm-AER2008 and Allen-Arkolakis-QJE2014.
To develop intuition for the proposed distinction between local and global dispersion forces, consider a hypothetical many-location economy with agents—such as consumers or firms—who make location choices. Suppose there are some agglomeration forces so that agents benefit from proximity to others. In the absence of any countervailing dispersion force, all agents concentrate in a single location (Fig. 1D).
Now, consider “local” dispersion forces: negative externalities that arise within a location, depend on its own population, and affect only its residents. A representative example is crowding in non-tradable sectors, such as housing scarcity. Specifically, if housing supply is inelastic, population growth within a city drives up prices, creating incentives for residents to relocate to nearby, less expensive areas. Agents relocate incrementally toward the periphery to mitigate congestion while still benefiting from proximity to others. In equilibrium, this tension gives rise to a single-peaked, monocentric spatial pattern, with a dominant center surrounded by lower-density fringes, as shown in Fig. 1E.
By contrast, “global” dispersion forces are defined as negative spillovers across locations. A canonical example is the market crowding effect resulting from interregional trade, combined with immobile factors such as land or other natural resources. In this case, interregional trade allows the central location with large agglomeration benefits to outcompete neighboring areas, leading to revenue losses in the local markets of those regions. If transport costs are sufficiently high, producers have an incentive to relocate to more distant areas to escape intense price competition at the center and to better capture local demand, while benefiting from local immobile resources such as land or natural endowments. This creates incentives for firms and workers to spread out across space, giving rise to multiple agglomerations that compete for market access and resources, and thereby spatially repel each other (Fig. 1F).
As transport costs decline, local and global dispersion forces exert countervailing influences. At the global scale, improved transport access extends the spatial reach of firms and consumers, intensifying market crowding. This undermines smaller agglomerations, driving economic activities to concentrate in a limited number of major hubs. At the local scale, however, lower transport costs reduce the relative advantages of proximity, fostering spatial spread within cities due to congestion forces. These opposing mechanisms can jointly give rise to a dual pattern: economy-wide concentration alongside intraurban decentralization.
The concepts of local and global dispersion forces have important implications for spatial policy analysis. This study shows that, while there are various possible microfoundations for these forces, distinguishing between them helps clarify the equilibrium structure and comparative statics of models of economic agglomeration. In particular, we provide a systematic taxonomy of spatial models into three broad model classes: (i) those with only local dispersion forces, (ii) those with only global dispersion forces, and (iii) those that incorporate both. These model classes can imply fundamentally different predictions regarding the spatial impact of declining transport costs. Thus, quantitative models can yield sharply divergent counterfactual predictions depending on the dispersion forces they embed. For example, declining transport costs tend to promote spatial spreading under local dispersion forces (e.g., Helpman-Book1998; Allen-Arkolakis-QJE2014), but foster further agglomeration under global dispersion forces (e.g., Krugman-JPE1991). Accordingly, transport policies intended to support peripheral regions may succeed or backfire, depending on whether the model used to design them adequately represents the dispersion forces relevant to the spatial problem at hand.
2   The two-region economy
We begin with a parsimonious two-region framework to introduce key concepts, and later discuss richer settings including general equilibrium models (e.g., Krugman-JPE1991; Helpman-Book1998; Tabuchi-JUE1998; Redding-Sturm-AER2008; Allen-Arkolakis-QJE2014). Then, Section 3 will consider a stylized many-region setting.
Consider a closed two-region economy, and let $\mathcal{I}\equiv\{1,2\}$ be the set of regions. There are a continuum of perfectly mobile agents, such as households, who choose their location to maximize utility. Let $x_{i}\geq 0$ denote the mass of agents in region $i$. The spatial distribution is $\bm{x}=(x_{1},x_{2})$. We let $x_{1}+x_{2}=1$ for normalization. Let $v_{i}(\bm{x})$ denote the indirect utility of agents in region $i$ as a function of $\bm{x}$. A spatial equilibrium is a distribution $\bm{x}$ such that no agent has an incentive to relocate. [^3] Formally, $\bm{x}$ is an equilibrium if $\bar{v}\geq v_{i}(\bm{x})$ for all $i$ with equality if $x_{i}>0$, where $\bar{v}$ is an endogenous scalar representing the corresponding equilibrium utility level.
Throughout this section, we consider a series of specifications for the indirect utility function $v_{i}(\bm{x})$ to highlight the difference between “local” and “global” dispersion forces. In doing so, we assume that regional characteristics are symmetric to isolate the trade-offs between agglomeration and dispersion forces driven by transport costs and the endogenous spatial distribution of mobile agents.
Transport across the regions is costly and captured by a proximity matrix $[\phi_{ij}]$, where $\phi_{ij}\in(0,1]$ measures the ease of transport from region $i$ to $j$. Higher values indicate better access. Throughout this section, we assume that
$\begin{bmatrix}\phi_{11}&\phi_{12}\\ \phi_{21}&\phi_{11}\end{bmatrix}=\begin{bmatrix}1&\phi\\ \phi&1\end{bmatrix},$ (1)
where $\phi\in(0,1)$ is the transport cost parameter.
2.1   The Beckmann model: A “local” dispersion force
We begin with a minimal model that exemplifies a “local” dispersion mechanism. This also serves to revisit the classic intuition of agglomeration as symmetry-breaking.
Consider the following reduced-form model featuring the trade-off between intraregional congestion costs and agents’ desire to be close to each other:
$v_{i}(\bm{x})=x_{i}^{-\beta}\Big{(}\sum_{j\in\mathcal{I}}\phi_{ij}x_{j}\Big{)}^{\alpha}$ $(\alpha,\beta>0).$ (2)
The first term ($x_{i}^{-\beta}$) captures congestion force, while the second reflects proximity-based positive externalities. We refer to this model as the Beckmann model after the seminal work by Beckmann-Book1976.444Beckmann studied spatial agglomeration in a continuous one-dimensional space. We consider its discrete-space and multiplicative version to streamline the exposition.
The uniform distribution of agents $\bar{\bm{x}}=(\tfrac{1}{2},\tfrac{1}{2})$ is an equilibrium, as both regions yield equal utility at $\bar{\bm{x}}$. A crucial feature of the model is that $\bar{\bm{x}}$ may not be stable. If the agglomeration force dominates the dispersion force, a tiny migration shock encourages further migration of agents, pushing the system away from $\bar{\bm{x}}$. Accordingly, a spatial pattern with $x_{1}>x_{2}$ or $x_{1}<x_{2}$ becomes a stable equilibrium. That is, agglomeration of mobile agents occurs without ex-ante regional asymmetries.
To determine the stability of $\bar{\bm{x}}$, we must check the incentive of a hypothetical migrant. Specifically, the utility gain from migration of a marginal mover from region $2$ to $1$ can be measured by the following elasticity of the utility difference:
$\omega$ $\equiv\frac{\bar{x}}{\bar{v}}\times\left.\frac{\partial{\left(v_{1}(\bm{x})-v_{2}(\bm{x})\right)}}{\partial{x_{1}}}\right|_{\bm{x}=\bar{\bm{x}}}=\frac{\bar{x}}{\bar{v}}\times\left.\left(\frac{\partial{v_{1}(\bm{x})}}{\partial{x_{1}}}-\frac{\partial{v_{2}(\bm{x})}}{\partial{x_{1}}}\right)\right|_{\bm{x}=\bar{\bm{x}}},$ (3)
where $\bar{x}\equiv\tfrac{1}{2}$ and $\bar{v}=v_{1}(\bar{\bm{x}})=v_{2}(\bar{\bm{x}})$ is the equilibrium utility level at $\bar{\bm{x}}$. The term $\frac{\bar{x}}{\bar{v}}$ simplifies the resulting expression.
If $\omega<0$, it indicates that a marginal increase in the mass of agents in region $1$ induces a relative utility decrease therein compared to region $2$. Migrants are strictly worse off, implying reverse migration. Hence, $\bar{\bm{x}}$ is stable. If $\omega>0$ instead, migrants are strictly better off. Further migration should occur. That is, $\bar{\bm{x}}$ is unstable.
Evaluating Eq. 3 for the Beckmann model yields
$\omega$ $=-\beta+\alpha\Theta,\quad\text{where}\quad\Theta\equiv\frac{1-\phi}{1+\phi}.$ (4)
The term $\Theta\in(0,1)$ is a convenient index of transport cost. It decreases as $\phi$ increases, with $\Theta\to 1$ as $\phi\to 0$ (regional autarky) and $\Theta\to 0$ as $\phi\to 1$ (frictionless transport).
The expression of $\omega$ is useful in two respects. First, it succinctly captures the trade-off between two opposing forces. The negative term $-\beta$ reflects the dispersion force, whereas the positive term $\alpha\Theta$ reflects the agglomeration force. The stability of $\bar{\bm{x}}$ (i.e., the sign of $\omega$) depends on which force dominates. Specifically,
$\omega\lesseqgtr 0\quad\Leftrightarrow\quad\alpha\Theta\lesseqgtr\beta$ (5)
with the same signs for the inequalities. For example, if congestion force is too strong ($\beta\geq\alpha$), $\omega<0$ for all $\phi\in(0,1)$, implying no agglomeration can occur.
Second, the expression also elucidates the spatial nature of the economic forces. The agglomeration force depends on transport costs through $\Theta$, whereas the dispersion force does not. The dispersion force is strictly “local” in that it is determined solely by the population within each region, is borne only by local residents, and remains unaffected by interregional transport conditions. By contrast, the agglomeration force is “global,” as it incorporates interregional interactions. As distance becomes less costly (as $\Theta$ decreases), the agglomeration force $\alpha\Theta$ weakens, since the benefit of being close to others diminishes. [^5] Note that, in the opposite extreme of $\phi\to 0$, $\Theta\to 1$ and $\alpha\Theta\to\alpha$. Since the regions are in autarky, the agglomeration force reduces to local spillovers à la Henderson-AER1974, and Eq. 5 reduces to whether positive spillovers $\alpha$ is smaller than negative ones $\beta$ or not. If $\beta<\alpha$, transport costs can matter: $\bar{\bm{x}}$ is stable for low transport costs (small $\Theta$Leftrightarrow large $\phi$) but unstable for high transport costs (large $\Theta$Leftrightarrow small $\phi$).
2.2   The Braid model: A “global” dispersion force
Next, we consider another minimal model with a “global” dispersion force. We replace the local congestion force in the Beckmann model with a trade-based dispersion force:
$v_{i}(\bm{x})=y_{i}(\bm{x})\Big{(}\sum_{j}\phi_{ij}x_{j}\Big{)}^{\alpha},$ (6)
where $y_{i}(\bm{x})$ is the income of an agent in region $i$ under given $\bm{x}$, which depends on costly transport. Specifically, the economy features gravity-type trade flows, whose microfoundations abstracted away for brevity. Each region $m$ is endowed with one unit of expenditure, which is allocated across all agents according to accessibility. The share received by an agent located in region $i$ from region $m$ is given by
$s_{i\mid m}\equiv\frac{\phi_{im}}{\sum_{l\in\mathcal{I}}x_{l}\phi_{lm}},$ (7)
ensuring that total spending from region $m$ equals one: $\sum_{i\in\mathcal{I}}x_{i}s_{i\mid m}=1$. The income of an agent in region $i$ is then given by
$y_{i}(\bm{x})=\sum_{m\in\mathcal{I}}s_{i\mid m}=\sum_{m\in\mathcal{I}}\frac{\phi_{im}}{\sum_{l\in\mathcal{I}}x_{l}\phi_{lm}}.$ (8)
For convenience, we refer this reduced-form model [^6] The model does not specify how consumers’ utility depends on their consumption of the tradables, where the expenditure comes from, what is the technology of producing tradables, how the prices of tradables are determined in equilibrium. as the Braid model after Braid-RSUE1988 who studied the case $\alpha=0$.
The uniform distribution is an equilibrium. Evaluating Eq. 3 gives
$\omega=\alpha\Theta-\Theta^{2}.$ (9)
Agglomeration occurs as long as $\alpha>0$. If $10<\alpha<$ 1, $\bar{\bm{x}}$ is stable for high transport costs (large $\Theta$Leftrightarrow small $\phi$) and unstable for low transport costs (small $\Theta$Leftrightarrow large $\phi$). [^7] If $\alpha\geq 1$, the symmetry is unstable for all $\phi\in(0,1)$, as $\Theta<1$ implies $\omega=(\alpha-\Theta)\Theta>0$.
Importantly, $\omega$ features a “global” dispersion force—a negative term in $\Theta$. The dispersion force stems from an interregional crowding effect embedded the share formula (Eq. 7). For an agent in region $i$, an increase in the mass of agents in any location implies less income from any other region $m$ because $s_{i\mid m}$ is smaller if region $m$ has better access to other locations (i.e., if $\sum_{l\in\mathcal{I}}x_{l}\phi_{lm}$ is larger). Concretely, we see
$\frac{\partial s_{i\mid m}}{\partial x_{j}}=-\frac{\phi_{im}}{\sum_{l\in\mathcal{I}}x_{l}\phi_{lm}}\times\frac{\phi_{jm}}{\sum_{l\in\mathcal{I}}x_{l}\phi_{lm}}=-s_{i\mid m}\times s_{j\mid m}<0$ (10)
for all $j$. The magnitude of this negative effect depends on both $\phi_{im}$ and $\phi_{jm}$. As such, this must appear as a second-order term of the transport cost index $\Theta$.
2.3   Contrasting implications of transportation costs
[…]
Population of region 1
A The Beckmann model (Eq. 2)
[…]
Population of region 1
B The Braid model (Eqs. 6 and 8)
[…]
Population of region 1
C The Beckmann model with asymmetry
[…]
Population of region 1
D The Braid model with asymmetry
Figure 3: Equilibrium values of $x_{1}$ in the Braid model and the Beckmann model
Note: We set $\alpha=1/2$ for both models, and $\beta=1/6$ for the Beckmann model. In Panels (A) and (B), the black markers indicate the points where the symmetry becomes unstable. In Panels (C) and (D), we multiply $v_{1}(\bm{x})$ by $1.05$. The black markers indicate the points at which stable and unstable equilibrium curves converge. For reference, the equilibria for the symmetric cases are shown in light gray.
Evidently, the Braid model and the Beckmann model offer contrasting implications regarding how transport costs shape endogenous asymmetries in the spatial distribution. Figures 3B and 3A highlight the contrast by showing the full equilibrium paths over the $\phi$-axis. In Fig. 3A, the symmetric equilibrium in the Beckmann model is stable if transport costs are low (i.e., if $\phi$ is large), but agglomeration emerges as transport costs increase (as $\phi$ decreases). The Braid model exhibits the opposite pattern, as shown in Fig. 3B.
These tendencies persist under the presence of asymmetries. In Figs. 3D and 3C, region 1 has an exogenous advantage. The symmetric pattern is not an equilibrium anymore. However, focusing on the equilibrium paths with $x_{1}>x_{2}$, broad implications are the same as the symmetric case. Lowering transport costs leads to dispersion in the Beckmann model, but to agglomeration in the Braid model.
While the literature has not explicitly referred to the spatial nature of dispersion forces, it is recognized that various dispersion mechanisms have opposing effects. In particular, Fujita-Thisse-Book2013, Chapter 8, compared the seminal models by Krugman-JPE1991 and Helpman-Book1998 and noted that “Krugman’s scenario is reversed” (p.289) in Helpman-like frameworks with urban costs. That is, agglomeration is preferred to dispersion if transport costs are sufficiently low in the former, while the converse is true for the latter. In fact, the minimal models capture the essence of these frameworks, as discussed in the remainder of this section.
2.4   Benefit matrix and the spatial scale of economic forces
The discussed contrast illustrates how the spatial scale of dispersion forces—local or global—can alter the spatial implications of declining transport costs. To formalize this insight and provide a tool for analyzing more sophisticated models, we now introduce the benefit matrix. For each specific indirect utility function $\bm{v}$, we define its benefit matrix as its elasticity matrix at the symmetric equilibrium, $\mathbf{V}\equiv\frac{\bar{x}}{\bar{v}}[\frac{\partial v_{i}}{\partial x_{j}}(\bar{\bm{x}})]_{i\in\mathcal{I},j\in\mathcal{I}}$. It summarizes how changes in the population distribution affect individual utility in the model under consideration. This can be formalized as follows:
Observation 1.
For each specification of $\bm{v}$, the utility gain $\omega$ at the symmetric equilibrium is an eigenvalue of the benefit matrix $\mathbf{V}$, and the corresponding eigenvector is $\bm{z}=(1,-1)$. ∎
[…]
$\mathbf{V}=-\beta\mathbf{I}+\alpha\mathbf{D}\quad\text{with}\quad\mathbf{D}=\frac{1}{1+\phi}\begin{bmatrix}1&\phi\\ \phi&1\end{bmatrix},$ (11)
where $\mathbf{I}$ is the identity matrix and $\mathbf{D}$ is the row-normalized proximity matrix. Observe that the expression for the utility gain, $\omega=-\beta+\alpha\Theta$, mirrors the structure of $\mathbf{V}$. In fact, the index $\Theta=\frac{1-\phi}{1+\phi}$ is an eigenvalue of $\mathbf{D}$, with the eigenvector being $\bm{z}$. It follows that $\mathbf{V}\bm{z}=(-\beta+\alpha\Theta)\bm{z}=\omega\bm{z}$, showing that $\omega$ is indeed an eigenvalue of $\mathbf{V}$.
Importantly, the representation of $\mathbf{V}$ based on $\mathbf{D}$ succinctly summarizes how agglomeration and dispersion forces interact with transport costs in the model at hand, as $\mathbf{D}$ completely contains the underlying transport cost structure.
In the case of the Beckmann model, the term $-\beta\mathbf{I}$ arises from the local congestion effects $x_{i}^{-\beta}$ and hence it is independent of $\mathbf{D}$. The term $\alpha\mathbf{D}$ reflects the proximity-based externalities and thus incorporates $\mathbf{D}$.
Similarly, for the Braid model, we can show that
$\mathbf{V}=\alpha\mathbf{D}-\mathbf{D}^{2}.$ (12)
This also implies $\omega=\alpha\Theta-\Theta^{2}$ as discussed earlier. The first term is the same agglomeration force as the Beckmann model. The second-order term $-\mathbf{D}^{2}$ captures the dispersion force stemming from competition across agents over dispersed demand.
These observations suggest a systematic way to distinguish the spatial scale of economic forces. Suppose the benefit matrix $\mathbf{V}$ is a polynomial in $\mathbf{D}$, such as $\mathbf{V}=c_{0}\mathbf{I}+c_{1}\mathbf{D}+c_{2}\mathbf{D}^{2}$. Then, we define an economic force as local if it appears as the coefficient of $\mathbf{I}$, and global if it arises from the terms involving powers of $\mathbf{D}$ [^9] A subtle point is that the distinction between local and global forces reflects modeling choices rather than universally applicable interpretations of the underlying economic mechanisms. For example, a “congestion force” may be a global force if it is formulated as crowding of regional facilities that are accessible from other regions through costly transport.
2.5   General equilibrium models
The benefit matrix approach is particularly effective for understanding general equilibrium models. For various models in the literature, the benefit matrix $\mathbf{V}$ can be cleanly represented by $\mathbf{D}$, which facilitates systematic comparison. Below, we consider a series of notable models assuming the symmetric two-region economy, and discuss them through the lens of local and global dispersion forces. Derivations are provided in Appendix F in the online appendix.
The Helpman (1998) / Redding–Sturm (2008) model. The Helpman-Book1998 model considers an imperfectly competitive framework with a horizontally differentiated sector for tradable goods, and introduces the consumption of local, non-tradable goods (i.e., housing). We consider its variant proposed by Redding-Sturm-AER2008.
For illustration, we provide a slightly detailed derivation for the benefit matrix $\mathbf{V}$ for this model. The indirect utility of mobile workers is given by
$v_{i}(\bm{x})=x_{i}^{-\gamma}w_{i}^{\mu}\cdot\textsc{CMA}_{i}^{\frac{\mu}{\sigma-1}},$ (13)
where $\mu\in(0,1)$ is consumers’ expenditure share on tradables and $\gamma\equiv 1-\mu$ is that on non-tradables, $\sigma>1$ is the elasticity of substitution of horizontally differentiated tradable varieties, $w_{i}$ is the wage in region $i$, and $\textsc{CMA}_{i}\equiv\sum_{j\in\mathcal{I}}x_{j}w_{j}^{1-\sigma}\phi_{ji}$ is the so-called “consumer market access.” The proximity matrix for this model is $\phi_{ij}=\tau_{ij}^{1-\sigma}$, where $\tau_{ij}\geq 1$ is the iceberg trade cost from region $i$ to $j$. Given $\bm{x}$, the wage $\bm{w}=(w_{i})$ is an endogenous object determined in general equilibrium with interregional trade.
The benefit matrix at $\bar{\bm{x}}$ reflects the structure of the indirect utility:
$\mathbf{V}$ $=-\gamma\mathbf{I}+\mu\mathbf{W}+\frac{\mu}{\sigma-1}\left(\mathbf{D}-(\sigma-1)\mathbf{D}\mathbf{W}\right).$ (14)
Here, $\mathbf{W}$ is the elasticity matrix for wages $\bm{w}$ at $\bar{\bm{x}}$, i.e., $\mathbf{W}\equiv\frac{\bar{x}}{\bar{w}}[\frac{\partial w_{i}}{\partial x_{j}}(\bar{\bm{x}})]_{i\in\mathcal{I},j\in\mathcal{I}}$ with $\bar{w}$ being the uniform level of nominal wages at $\bar{\bm{x}}$. It represents the general equilibrium response of wages to marginal migration, and given explicitly as
$\mathbf{W}=\frac{1}{\sigma}C(\mathbf{D})\cdot\mathbf{D},\quad\text{where}\quad C(\mathbf{D})\equiv\left(\mathbf{I}+\frac{\sigma-1}{\sigma}\mathbf{D}\right)^{-1}.$ (15)
In Eq. 14, the first component of $\mathbf{V}$ is a local dispersion force arising from crowding in housing markets. The second term is the response of wages. The third component is the so-called “price index effects,” the global forces driven by consumer market access. That is, the cost of living in a region falls when nearby regions offer a greater variety of goods, but rises when they pay higher wages, thereby pushing up local prices.
Substituting Eq. 15 into Eq. 14 and rearranging, $\mathbf{V}$ reduces to
$\mathbf{V}=C(\mathbf{D})\cdot\left(-\gamma\mathbf{I}+c_{1}\mathbf{D}\right),$ (16)
[…]
The final expression of $\omega$ includes a complicating term $C(\Theta)$ stemming from general equilibrium. However, since $C(\Theta)>0$, the stability of $\bar{\bm{x}}$ hinges on the sign of the linear expression $-\gamma+c_{1}\Theta$. Observe that there is a local dispersion force $-\gamma$. Also, $c_{1}$ encapsulates both positive and negative effects concerning spatial agglomeration. If $\mu$ is sufficiently large, then $c_{1}>0$ and there is a (net) global agglomeration force. [^11] In general, the sign of $c_{1}$ is indeterminate as there are both agglomeration and dispersion forces. Each term in $c_{1}$ has a clear economic interpretation as extensively discussed elsewhere (see, e.g., Fujita-Thisse-Book2013; Baldwin-et-al-Book2003; Brakman-etal-Book2019, for surveys). Since model-specific interpretations are not our main focus, such discussions are relegated to Appendix F in the online appendix. If $\sigma(1-\mu)<1$, $\bar{\bm{x}}$ is unstable for high transport costs and stable for low transport costs.
Although the key coefficients $-\gamma$ and $c_{1}$ are microfounded through general equilibrium, the implication on the stability of $\bar{\bm{x}}$ resembles that of the Beckmann model. The central dispersion mechanism in the Helpman framework is crowding in non-tradables markets, whose archetype is local congestion as in the Beckmann model.
The Allen–Arkolakis (2014) model. Allen-Arkolakis-QJE2014 proposed a perfectly competitive framework with externalities in the form of within-region congestion and positive productivity spillovers. As shown in the original paper, this model bears a structural similarity to the Helpman framework. In fact, we can show that
$\mathbf{V}=C(\mathbf{D})\cdot\left(c_{0}\mathbf{I}+c_{1}\mathbf{D}\right)$ (17)
where $C(\mathbf{D})$ captures general equilibrium adjustments, and we set $c_{0}\equiv\alpha+\beta-\frac{1+\alpha}{\sigma}$ and $c_{1}\equiv\alpha+\beta+\frac{1-\beta}{\sigma}$. Here, $\alpha>0$ and $\beta<0$ denote the elasticities of local agglomeration and congestion effects with respect to local population, respectively. As in the Helpman model, the sign of the utility gain $\omega$ again depends on a linear expression $c_{0}+c_{1}\Theta$. If agglomeration forces are sufficiently strong, the model can exhibit agglomeration under high transport costs and dispersion under low transport costs. [^12] As in the Beckmann model, sufficiently strong local dispersion force suppresses agglomeration: $|\beta|>\alpha$ implies $\omega<0$ for all transport cost levels. As in the Braid model, the symmetry is unstable for all transport cost levels if agglomeration force $\alpha$ is too large. We can thus view the Helpman model and the Allen–Arkolakis model as models driven by local dispersion forces.
The Krugman (1991) model. On the other hand, Krugman-JPE1991 emphasizes the role of market crowding, a global dispersion force, by assuming immobile consumers dispersed across regions. For this model, we compute that
$\mathbf{V}=\mathbf{W}+\frac{\mu}{\sigma-1}\left(\mathbf{D}-(\sigma-1)\mathbf{D}\mathbf{W}\right).$ (18)
Naturally, local dispersion forces are absent in $\mathbf{V}$. Furthermore, in this model,
$\mathbf{W}=\frac{1}{\sigma}C(\mathbf{D})\cdot\left(\mu\mathbf{D}-\mathbf{D}^{2}\right),\quad\text{where}\quad C(\mathbf{D})\equiv\left(\mathbf{I}-\frac{\mu}{\sigma}\mathbf{D}-\frac{\sigma-1}{\sigma}\mathbf{D}^{2}\right)^{-1}.$ (19)
Comparing $\mathbf{W}$ in Eqs. 15 and 19, we observe that the Krugman model has a negative term $-\mathbf{D}^{2}$, while the Helpman model does not. The benefit matrix is given by
$\mathbf{V}=C(\mathbf{D})\cdot\left(c_{1}\mathbf{D}-c_{2}\mathbf{D}^{2}\right).$ (20)
where $c_{1}\equiv\frac{\mu}{\sigma-1}+\frac{\mu}{\sigma}>0$ and $c_{2}\equiv\frac{\mu^{2}}{\sigma-1}+\frac{1}{\sigma}>0$. This implies $\omega=C(\Theta)\cdot(c_{1}\Theta-c_{2}\Theta^{2})$ with $C(\Theta)\equiv 1/(1-\frac{\mu}{\sigma}\Theta-\frac{\sigma-1}{\sigma}\Theta^{2})>0$. Thus, the core trade-off between agglomeration and dispersion forces in the model is encapsulated by a concave quadratic; $c_{1}\Theta>0$ represents global agglomeration forces, and $-c_{2}\Theta^{2}<0$ captures global dispersion forces augmented by immobile demands. In this model, there is always a nonzero demand for the manufacturing goods in each region due to the presence of immobile consumers. This discourages concentration of production in one region if transport costs are high, and the uniform distribution becomes stable. As in the Braid model, agglomeration occurs if transport costs decrease from a high level.
The Tabuchi (1998) model. Tabuchi-JUE1998 extends Krugman’s model by introducing urban costs while retaining immobile consumers. In addition to the regional-scale component of Krugman, each region has an Alonso–Muth–Mills monocentric city structure. Within each region, agents commute to a central business district and face the trade-off between commuting costs and land rent. The benefit matrix reduces to
$\mathbf{V}=C(\mathbf{D})\cdot\left(-c_{0}\mathbf{I}+c_{1}\mathbf{D}-c_{2}\mathbf{D}^{2}\right),$ (21)
where $C(\mathbf{D})$ is a matrix factor capturing general equilibrium effects, and $c_{0}=\gamma\epsilon_{1}>0$, $c_{1}=\frac{\mu}{\sigma-1}+\frac{\mu}{\sigma}>0$, and $c_{2}=\frac{\mu^{2}}{\sigma-1}\epsilon_{2}+\frac{1}{\sigma}\epsilon_{3}$. Here, $\gamma$ is the expenditure share on housing, and $\epsilon_{1},\epsilon_{2},\epsilon_{3}$ capture the effects of urban costs. For example, the larger the intraregional commuting cost parameter and/or the opportunity cost of land, the larger $\epsilon_{1}>0$ becomes, so that local dispersion force $-c_{0}$ becomes greater in magnitude. The stability of $\bar{\bm{x}}$ depends on a quadratic expression $-c_{0}+c_{1}\Theta-c_{2}\Theta^{2}$.
The agglomeration force ($c_{1}\Theta>0$) is the same as the Krugman model, while the local dispersion force ($-c_{0}<0$) is due to urban costs. The term $c_{2}\Theta^{2}$ captures global forces arising from interregional trade as well as the impacts of urban costs. The model features both local and global dispersion forces. The symmetry can be stable for both high and low levels of transport costs.
Idiosyncratic taste shocks. It is worthwhile to mention idiosyncratic preference shocks (McFadden-Book1974; McFadden-Book1978; McFadden-TRB1978) often considered in both quantitative and theoretical literature (Hunt-Simmonds-EPB1993; Waddell-JAP2002; Anas-Liu-JRS2007; Redding-Rossi-Hansberg-ARE2017; anderson1992discrete; Tabuchi-Thisse-JDE2002; Murata-JUE2003). On such taste shocks, Behrens-Murata-DP2018 demonstrated that the equilibrium condition for spatial distributions in models with idiosyncratic shocks can be equivalently represented by that in homogeneous preference models with local non-tradables markets as in Helpman. Thus, from the perspective of stabilizing forces at the symmetry, introducing idiosyncratic shocks into a model with homogeneous preference is similar to adding a negative constant to the utility gain $\omega$, i.e., to embedding an extra local dispersion force. Welfare implications are, however, different and care should be taken (Behrens-Murata-DP2018).
2.6   The three model classes
As the preceding models illustrate, the central trade-offs between agglomeration and dispersion forces often take a quadratic form in $\mathbf{D}$. That is, for all specifications of the indirect utility function $\bm{v}$ considered so far, the benefit matrix takes the form $\mathbf{V}=\mathbf{C}\cdot\big{(}c_{0}\mathbf{I}+c_{1}\mathbf{D}+c_{2}\mathbf{D}^{2}\big{)}$, with $\mathbf{C}$ and $\{c_{0},c_{1},c_{2}\}$ being model-dependent. Further, only the quadratic component determines the stability of $\bar{\bm{x}}$ in the sense that $\operatorname{sgn}\left[\omega\right]=\operatorname{sgn}\left[c_{0}+c_{1}\Theta+c_{2}\Theta^{2}\right]$ where $\operatorname{sgn}[\cdot]$ denote the sign. For brevity, we denote
$\mathbf{V}\propto c_{0}\mathbf{I}+c_{1}\mathbf{D}+c_{2}\mathbf{D}^{2},$ (22)
to indicate that such conditions are satisfied for the benefit matrix $\mathbf{V}$.
[…]
A Allen–Arkolakis (Type L)
[…]
B Krugman (Type G)
[…]
C Tabuchi (Type LG)
Figure 4: Representative forms of the quadratic component of the utility gain
Note: For each selected example from the three model types, the relevant quadratic component of the utility gain is plotted as a function of the transport cost index $\Theta$. The parameter values are chosen such that the stability of symmetry switches as $\Theta$ varies. In each panel, the symmetric pattern is stable for $\Theta$ such that the curve lies below the horizontal axis. In Panels (A) and (C), the intersept at $\Theta=0$ corresponds to the local dispersion force $c_{0}<0$. Local dispersion forces stabilize $\bar{\bm{x}}$ under low transport costs, and global dispersion forces under high transport costs.
Model Class Stability of symmetry Spatial scale of dispersion force Examples Type L Low transport costs Local Beckmann-Book1976 Helpman-Book1998 Murata-Thisse-JUE2005 Redding-Sturm-AER2008 Allen-Arkolakis-QJE2014 Redding-Rossi-Hansberg-ARE2017, Section 3 Type G High transport costs Global Harris-Wilson-EPA1978 Krugman-JPE1991 Krugman-Venables-QJE1995 Puga-EER1999, Section 3 Forslid-Ottaviano-JoEG2003 Pfluger-RSUE2003 Type LG High and low transport costs Both local and global Tabuchi-JUE1998 Fujita-Krugman-Venables-Book1999, Section 14.4 Puga-EER1999, Section 4 Pfluger-Suedekum-JUE2008 Pfluger-Tabuchi-RSUE2010 Kucheryavyy-etal-JIE2024
Table 1: Notable examples for the three model classes
Note: The section numbers in the table corresponds to those in the referenced papers. For Krugman-Venables-QJE1995 and Puga-EER1999, the share of manufacturing sector is interpreted as the spatial distribution of interest. In all the models in the table, the stability of $\bar{\bm{x}}$ hinges on the sign of a quadratic function of the form $c_{0}+c_{1}\Theta+c_{2}\Theta^{2}$ with model-dependent coefficients $\{c_{0},c_{1},c_{2}\}$. For Type L models, $c_{0}<0$ and $c_{0}+c_{1}+c_{2}>0$, so that $\bar{\bm{x}}$ is unstable for large $\Theta$. Some Type L models incorporate market crowding effects across regions, meaning that $c_{2}$ can be negative. However, for any values of structural parameters that allow for multiple equilibria, the symmetry is unstable when $\Theta$ is large. For Type G models, $c_{0}\geq 0$, indicating that there is no net local dispersion forces.
Under the quadratic representation in Eq. 22, we can define the three prototypical model classes mentioned in Section 1. Specifically, we classify models depending on the conditions under which symmetry can be stable: high transport costs (large $\Theta$), low transport costs (small $\Theta$), or both. We refer to the three model classes as Type L, Type G, and Type LG, where L denotes “local” and G denotes “global.” We can summarize as follows, with $\{c_{k}\}$ being model-dependent constants: [^13] See Section A.2 for the formal definitions. To classify models, we assume that agglomeration forces are sufficiently strong so that the stability of the symmetric equilibrium depends on the transport cost index $\Theta$, meaning that $\omega$ cuts the horizontal axis at some $\Theta$. This excludes the case in which the symmetry is stable (or unstable) for all $\phi$.
  • Type L emphasizes local dispersion forces. The symmetry is stable if transport costs are low, and unstable if transport costs are high. A typical form is $\mathbf{V}\propto c_{0}\mathbf{I}+c_{1}\mathbf{D}$ with $c_{0}<0$ indicating the local dispersion force.
  • Type G emphasizes global dispersion forces. The symmetry can be stable if transport costs are high, and unstable if transport costs are low. Typically, $\mathbf{V}\propto c_{1}\mathbf{D}+c_{2}\mathbf{D}^{2}$, with $c_{2}<0$ representing the global dispersion force.
  • Type LG is the synthesis of Type G and Type L, and has both local and global dispersion forces: $\mathbf{V}\propto c_{0}\mathbf{I}+c_{1}\mathbf{D}+c_{2}\mathbf{D}^{2}$, where $c_{0},c_{2}<0$ and $c_{1}>0$. The symmetry can become stable for high and low transport cost levels.
Figure 4 illustrates typical forms of the relevant quadratic component of the utility gain for the three model classes. Table 1 lists notable examples in the literature.
In the quadratic representation in Eq. 22, the coefficients $\{c_{0},c_{1},c_{2}\}$ are functions of the model’s structural parameters, excluding transport costs encapsulated in $\mathbf{D}$. The constant $c_{0}$ captures all the forces within each location, i.e., the local forces. The latter two coefficients $c_{1}$ and $c_{2}$ represent the global forces across locations, with different orders of $\mathbf{D}$. The first-order coefficient $c_{1}$ reflects the direct effects from one region to another, such as the agglomeration externalities in the Beckmann model. Similarly, $c_{2}$ captures indirect spillovers, such as the market crowding effect at intermediate regions, as seen in gravity-type structure of interregional trade.
Each coefficient summarizes the composite effect at the corresponding order of $\mathbf{D}$, and its sign indicates the net impact on $\omega$. For instance, suppose that there are both local agglomeration economies and diseconomies. Then, $c_{0}>0$ indicates that local agglomeration economies dominate diseconomies, and vice versa.
3   Many regions
Our analysis so far has highlighted that different dispersion forces become consequential at distinct transport cost regimes. Although this insight has been recognized in various forms in the theoretical literature, its implications on spatial patterns for settings with more than two regions remain underexplored. This section considers an $N$-region economy. All variables and functions such as $\bm{x}$ or $\bm{v}$ are straightforwardly extended.
We must define the “fully dispersed” state to study spatial agglomeration. A natural candidate in an $N$-region setting is the uniform spatial distribution $\bar{\bm{x}}\equiv(\frac{1}{N},\frac{1}{N},\ldots,\frac{1}{N})$. To employ this distribution as the analytical ground, we assume a symmetric geography following krugman1993number, in which the regions are homogeneous and equidistantly placed over a circle, and transport is possible only along the circumference:
Figure 5: $N$-region symmetric circle.
Assumption C.
The proximity matrix is given by $\phi_{ij}=\phi^{\ell_{ij}}$, where $\phi\in(0,1)$ is the ease of transport between two consecutive regions, and $\ell_{ij}\equiv\min\{|i-j|,N-|i-j|\}$ is the distance between regions $i$ and $j$ over the circumference. All regions are symmetric regarding their local fundamentals (e.g., innate amenity or productivity). ∎
Figure 5 illustrates the geography. Since every region has the same level of accessibility to the other regions in a circle, Assumption C abstracts away geographical advantages (Matsuyama-RIE2017), the advantages from each region’s unique position within transport networks. Combined with the perfect symmetry in other regional fundamentals, [^14] See Assumption S in Section A.2 for the formal version of this part of Assumption C. the uniform distribution $\bar{\bm{x}}$ is always a spatial equilibrium.
3.1   Benefit matrix and the stability of symmetry
If the symmetric equilibrium does not withstand small migration shocks, it is unstable and some form of agglomeration should occur. A migration shock can be represented as $\bar{\bm{x}}+\bm{z}$, where $\bm{z}=(z_{i})_{i\in\mathcal{I}}$ is a deviation pattern such that $\sum_{i\in\mathcal{I}}z_{i}=0$. For each region $i$, $z_{i}>0$ represents a population increase, and $z_{i}<0$ a decrease. Clearly, the $N$-region setting allows for a wider range of deviation patterns than the two-region setting. [^15] Specifically, the only deviation pattern allowed in the two-region case is $\bm{z}=(\epsilon,-\epsilon)$ with small $\epsilon$.
To assess the local stability of the full dispersion $\bar{\bm{x}}$, we examine whether any infinitesimal deviation pattern can improve the utility of movers. The following lemma shows that the benefit matrix $\mathbf{V}$ once again plays a central role:
Lemma 1.
Let $\omega^{*}$ be the maximum possible utility gain from migration for a relocating agent upon an infinitesimal population deviation at $\bar{\bm{x}}$. Then, $\omega^{*}$ equals the largest eigenvalue of the benefit matrix $\mathbf{V}=\frac{\bar{x}}{\bar{v}}[\frac{\partial{v_{i}}}{\partial{x_{j}}}(\bar{\bm{x}})]_{i,j\in\mathcal{I}}$, and $\omega^{*}$ is achieved by deviation toward the associated eigenvector direction $\bm{z}^{*}$. ∎
Proof.
See Section A.1. ∎
The stability of $\bar{\bm{x}}$ depends on the sign of $\omega^{*}$. If $\omega^{*}<0$, any deviation lowers the utility of movers, implying local stability. Conversely, if $\omega^{*}>0$, movers benefit from deviating, triggering further migration. As a result, $\bar{\bm{x}}$ becomes unstable, and population starts to agglomerate along the direction $\bm{z}^{*}$.
Thus, we need to understand the eigenvalues of the benefit matrix $\mathbf{V}$. Our setting allows a key simplification on this regard:
Observation 2.
Under Assumption C, for each model (i.e., for each specific indirect utility function $\bm{v}$), the structure of the benefit matrix $\mathbf{V}$ remains unchanged from its two-region version. Specifically, if $\mathbf{V}=\Omega(\mathbf{D})$ with some matrix function $\Omega$ in the symmetric two-region setting, then we can just replace $\mathbf{D}$ with its many-region version to obtain the many-region version of $\mathbf{V}$. ∎
This allows a transparent approach for the stability analysis of $\bar{\bm{x}}$. For example, the formula $\mathbf{V}=-\beta\mathbf{I}+\alpha\mathbf{D}$ for the Beckmann model remains valid in the $N$-region economy, provided that $\mathbf{I}$ and $\mathbf{D}$ are replaced with their $N$-region counterparts. Then, the $k$th eigenvalue of $\mathbf{V}$ is $\omega_{k}=-\beta+\alpha\Theta_{k}$, where $\Theta_{k}$ denotes the $k$th eigenvalue of $\mathbf{D}$ [^16] If $\bm{z}_{k}$ denotes the $k$th eigenvector of $\mathbf{D}$, $\mathbf{D}\bm{z}_{k}=\Theta_{k}\bm{z}_{k}$ implies $\mathbf{V}\bm{z}_{k}=\omega_{k}\bm{z}_{k}$. Here, each $\omega_{k}$ measures the hypothetical utility gain for migrants if deviation occurs toward the $k$th eigenvector direction. In turn, $\omega^{*}=\max_{k}\{\omega_{k}\}$, and the stability of $\bar{\bm{x}}$ can be easily determined even in the many-region case. Observation 2 implies that such a clean approach is possible for all the models discussed earlier.
3.2   Maximum and minimum gains in proximity
It is thus important to understand the eigenvalue–eigenvector pairs of $\mathbf{D}$ before considering specific models. Once they are obtained, $\{\omega_{k}\}$ and $\omega^{*}$ are readily available. In fact, the eigenvalues $\{\Theta_{k}\}$ of $\mathbf{D}$ has a clear interpretation:
Observation 3.
Each eigenvalue $\Theta_{k}$ of $\mathbf{D}$ measures the gain in proximity experienced by migrants when migration shocks occur in the direction of the corresponding eigenvector $\bm{z}_{k}$. ∎
To see this, consider a special case of the Beckmann model with $\beta=0$ and $\alpha=1$ where $\mathbf{V}=\mathbf{D}$ and $\omega_{k}=\Theta_{k}$. On the other hand, the indirect utility in this case is nothing but a parsimonious proximity measure $v_{i}(\bm{x})=\sum_{j\in\mathcal{I}}\phi_{ij}x_{j}$. This means that, each $\Theta_{k}$ ($=\omega_{k}$) represents the proximity gain for relocated agents when migration shock occurs toward the $\bm{z}_{k}$ direction.
As a concrete example, suppose $N=4$. The normalized proximity matrix is then
$\mathbf{D}=\frac{1}{1+2\phi+\phi^{2}}\begin{bmatrix}1&\phi&\phi^{2}&\phi\\ \phi&1&\phi&\phi^{2}\\ \phi^{2}&\phi&1&\phi\\ \phi&\phi^{2}&\phi&1\end{bmatrix}.$ (23)
There are two relevant eigenvalues: $\Theta_{1}=\frac{1-\phi}{1+\phi}$ and $\Theta_{2}=\big{(}\frac{1-\phi}{1+\phi}\big{)}^{2}$. An eigenvector associated with $\Theta_{1}$ is $\bm{z}_{1}=(1,0,-1,0)$, and represents a monocentric spatial pattern such as $\bar{\bm{x}}+\epsilon\bm{z}_{1}=(\bar{x},\bar{x}+\epsilon,\bar{x},\bar{x}-\epsilon)$. One region grows at the expense of another that is two steps away, creating a single center of attraction. In contrast, $\bm{z}_{2}=(1,-1,1,-1)$ associated with $\Theta_{2}$ corresponds to a polycentric agglomeration pattern $\bar{\bm{x}}+\epsilon\bm{z}_{2}=(\bar{x}+\epsilon,\bar{x}-\epsilon,\bar{x}+\epsilon,\bar{x}-\epsilon)$. Two regions located two steps apart grow symmetrically, while the two intermediate regions shrink.
There are three intuitive properties about $\Theta_{1}$ and $\Theta_{2}$. First, both are positive. Any deviation from the full dispersion induces some form of agglomeration, and thus increases the average proximity among movers. Second, each $\Theta_{k}$ decreases as $\phi$ increases. That is, the gain from proximity decreases if transport costs matter less. Third, $\Theta_{1}>\Theta_{2}$ at any value of $\phi$. Naturally, deviation toward a monocentric pattern induces a greater proximity gain than toward a polycentric pattern.
[…]
A $k=1$
B $k=2$
C $k=4$
D $k=8$
Figure 6: Examples of spatial patterns generated by $\bm{z}_{k}$ ($N=16$).
In fact, these properties generalize to the $N$-region case. [^17] Akamatsu-Takayama-Ikeda-JEDC2012, Lemma 4.2, provides the analytical formulae for $\{\Theta_{k}\}$ and $\{\bm{z}_{k}\}$, whereas the interpretation of $\{\Theta_{k}\}$ as proximity gains is newly given in the present study. Lemma 5 in Appendix A reproduces the relevant part of the aforementioned lemma, adapted to our context. In this case, each $\bm{z}_{k}$ corresponds to a deviation pattern with $k$ symmetric peaks, where $k=1,2,\ldots,\frac{N}{2}$, as illustrated in Fig. 6 for the $N=16$ case. Most importantly, the maximum and minimum gains in proximity are unambiguously determined:
Lemma 2.
Under Assumption C, $\max_{k}\Theta_{k}=\Theta_{1}$ and $\min_{k}\Theta_{k}=\Theta_{\frac{N}{2}}$ for all $\phi$. ∎
Proof.
See Akamatsu-Takayama-Ikeda-JEDC2012, Lemma 4.2. ∎
Intuitively, at any value of $\phi$, the monocentric agglomeration pattern ($k=1$) yields the largest proximity gain. The smallest gain arises from the pattern with the largest number of symmetric peaks, in which agents concentrate in every other region ($k=\frac{N}{2}$).
3.3   Contrasting implications for spatial patterns
Lemma 2 allows us to study the stability of the symmetric equilibrium, and to reveal the contrasting impact of local and global dispersion forces on spatial patterns.
In models with a local dispersion force, what matters is the maximum eigenvalue of $\mathbf{D}$, i.e., $\Theta_{1}$. For instance, in the Beckmann model, since $\omega_{k}=-\beta+\Theta_{k}$,
$\omega^{*}\gtreqless 0\quad\Leftrightarrow\quad\max_{k}\left\{-\beta+\alpha\Theta_{k}\right\}=-\beta+\alpha\max_{k}\left\{\Theta_{k}\right\}\gtreqless 0\quad\Leftrightarrow\quad\Theta_{1}\gtreqless\frac{\alpha}{\beta}.$ (24)
Figure 7A illustrates this for the $N=8$ case. The symmetry is stable if $\Theta_{1}$ is sufficiently small, i.e., if $\phi$ is sufficiently large. At the point where $\bar{\bm{x}}$ becomes unstable, a deviation of the form $\bar{\bm{x}}+\epsilon\bm{z}_{1}$ is self-reinforcing, and a single-peaked pattern as illustrated in Fig. 6A is favored by migrants.
[…]
A The Beckmann model (Type L)
[…]
B The Braid model (Type G)
Figure 7: Curves of $\omega_{k}$ for the two minimal models ($N=8$, $k=1,2,3,4$).
Note: The uniform distribution is stable if all curves $\{\omega_{k}\}_{k=1}^{4}$ stay below the horizontal axis (gray region). It becomes unstable at $\phi^{*}$, at which the largest eigenvalue cuts the axis. In the Beckmann model, $\omega_{1}$ is the first to cross the axis, whereas in the Braid model, it is $\omega_{4}$.
In models with a global dispersion force, the minimum eigenvalue of $\mathbf{D}$ is critical. For example, in the Braid model, $\omega_{k}=\alpha\Theta_{k}-\Theta_{k}^{2}=\Theta_{k}(\alpha-\Theta_{k})$. Since $\Theta_{k}>0$, we see
$\omega^{*}\gtreqless 0\quad\Leftrightarrow\quad\max_{k}\left\{\alpha-\Theta_{k}\right\}=\alpha-\min_{k}\left\{\Theta_{k}\right\}\gtreqless 0\quad\Leftrightarrow\quad\alpha\gtreqless\Theta_{\frac{N}{2}},$ (25)
Figure 7B illustrates this for the $N=8$ case. The symmetry is stable if $\Theta_{\frac{N}{2}}$ is sufficiently large, that is, if $\phi$ is sufficiently small. At the point where $\bar{\bm{x}}$ becomes unstable, the $\frac{N}{2}$-centric deviation pattern $\bm{z}_{\frac{N}{2}}$ attracts agents, leading to the emergence of a polycentric pattern with the largest number of symmetric economic centers (Fig. 6D).
The minimal examples demonstrate that the preferred spatial pattern at the onset of instability of $\bar{\bm{x}}$ is formally different depending on the spatial scale of the dispersion force in a model. In particular, a polycentric spatial pattern is preferred only under a global dispersion force. This insight from the reduced-form models can be extended to cover general equilibrium models, such as those discussed in Section 2.5. We can summarize as follows:
Proposition 1.
Consider a spatial model under Assumption C.
  1. (a)
    Consider a model of Type L or LG with local dispersion forces. Then, the uniform distribution $\bar{\bm{x}}$ can be stable at low transport cost levels. Suppose that the model parameters are set in a way that $\bar{\bm{x}}$ is stable for all $\phi\in(\phi^{*},1)$ with some $\phi^{*}\in(0,1)$, and becomes unstable at $\phi^{*}$, i.e., $\bar{\bm{x}}$ is unstable for $\phi$ slightly smaller than $\phi^{*}$. Then, a monocentric, single-peaked spatial equilibrium branches from $\bar{\bm{x}}$ at $\phi^{*}$.
  2. (b)
    Consider a model of Type G or LG with global dispersion forces. Then, $\bar{\bm{x}}$ can be stable at high transport cost levels. Suppose that the model parameters are set in a way that $\bar{\bm{x}}$ is stable for all $\phi\in(0,\phi^{*})$ with some $\phi^{*}\in(0,1)$, and becomes unstable at $\phi^{*}$, i.e., $\bar{\bm{x}}$ is unstable for $\phi$ slightly larger than $\phi^{*}$. Then, a polycentric spatial equilibrium with $\frac{N}{2}$ symmetric peaks branches from $\bar{\bm{x}}$ at $\phi^{*}$.
Proof.
See Section A.2. ∎
Note that Proposition 1 focuses on the case where (i) multiple equilibria are possible but at the same time (ii) $\bar{\bm{x}}$ is stable for some $\phi$. If local dispersion forces are sufficiently strong in a Type L (or LG) model, $\bar{\bm{x}}$ is stable for all $\phi\in(0,1)$, and $\phi^{*}$ mentioned in Proposition 1 (b) may not exist. In this case, the equilibrium is unique for all transport cost levels. This convenient property has allowed unambiguous counterfactual analysis in conventional quantitative spatial models. It is noted that, however, uniqueness of equilibrium implies that $\bar{\bm{x}}$ is the only possible outcome under the absence of exogenous geographical asymmetries. Also, for Type G models, $\bar{\bm{x}}$ can be unstable for all $\phi$ if agglomeration force is too strong. In this case, all agents may concentrate in one location. Proposition 1 excludes these cases so that the level of transport costs matters.
3.4   Evolution of spatial patterns
Another important caveat about Proposition 1 is that it relies on local stability analysis at $\bar{\bm{x}}$, and hence does not tell whether the instability of $\bar{\bm{x}}$ toward a polycentric deviation pattern indeed implies the existence of stable polycentric spatial equilibria in the vicinity of $\phi^{*}$. In fact, to study the path of stable spatial equilibria along the transport cost axis beyond Proposition 1, one must assume a specific indirect utility function, and thus possible characterizations are inherently model dependent. [^18] For example, Kucheryavyy-etal-JIE2024 focused on a specific but flexible two-region model that encompass Allen-Arkolakis-QJE2014 and Krugman-JPE1991 as special cases. They essentially showed that the agricultural sector à la Krugman-JPE1991 produces a global dispersion force that stabilize symmetry at high transport costs, which is consistent with our results. On the other hand, existence results for equilibria are available in general symmetric geographies (Ikeda-Murota-Book2014).
A The Helpman model (Type L)
B The Forslid–Ottaviano model (Type G)
Figure 8: Numerical illustration of Propositions 3 and 2
Nonetheless, there is indeed a sharp difference in multi-region agglomeration process under the influence of different dispersion mechanisms. In particular, polycentric spatial patterns can become stable only under a pronounced global dispersion force, which is consistent with Proposition 1. This is illustrated by the following formal results for selected models from Types L and G under Assumption C with $N=4$:
Proposition 2.
Consider the Type L model by Helpman-Book1998 in the four-region symmetric circle. Assume that multiple equilibria exists, and that the full dispersion $\bar{\bm{x}}=(\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4})$ is stable for large $\phi$. Then, there exists $\phi^{*}\in(0,1)$ such that $\bar{\bm{x}}$ is unstable for all $\phi\in(0,\phi^{*})$. Duocentric spatial equilibria of the form $(m_{1},m_{2},m_{1},m_{2})$ with $m_{1}>m_{2}$, if exist, are unstable for all $\phi\in(0,\phi^{*})$, implying that stable equilibria must be single-peaked for this range. The full dispersion $\bar{\bm{x}}$ is stable for all $\phi\in(\phi^{*},1)$, and a monocentric equilibrium path is connected to $\bar{\bm{x}}$ at $\phi^{*}$ [Proposition 1 (a)].
Proposition 3.
Consider the Type G model by Forslid-Ottaviano-JoEG2003 in the four-region symmetric circle. Assume that the full dispersion $\bar{\bm{x}}=(\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4})$ is stable for small $\phi$. Suppose that the initial state is $\bar{\bm{x}}$, and $\phi$ increases monotonically from 00. At some $\phi^{*}\in(0,1)$, $\bar{\bm{x}}$ becomes unstable as the duocentric deviation pattern $\bm{z}=(1,-1,1,-1)$ generates positive utility gain [Proposition 1 (b)]. The new stable equilibrium is the duocentric concentration, e.g., $(\frac{1}{2},0,\frac{1}{2},0)$. At some $\phi^{**}>\phi^{*}$, the duocentric equilibrium becomes unstable, as the deviation pattern $\bm{z}=(1,0,-1,0)$ now generates a positive utility gain. Finally, the full concentration in a single region such as $(1,0,0,0)$ becomes the stable spatial equilibrium for large $\phi$.
Proof of Propositions 3 and 2.
See AMT-2016. ∎
Figure 8 numerically illustrates Propositions 3 and 2, with schematic diagrams above showing the corresponding spatial patterns along the circular space. In particular, in the Helpman model, a monocentric distribution of the form $\bm{x}=(m_{2},m_{1},m_{2},m_{3})$ with $m_{1}>m_{2}>m_{3}$ is the stable equilibrium for $\phi\in(0,\phi^{*})$. As $\phi$ increases, it converges to the full dispersion, which is consistent with Proposition 1 (a).
As a further numerical example, Figure 9 depicts the typical evolution of spatial patterns, assuming $N=16$. We consider a monotonic decrease in transport costs (a monotonic increase in $\phi$), and follow a path of stable spatial equilibria.
[…]
A Type L
B Type G
C Type LG
Figure 9: Stable spatial patterns at different transport cost levels.
Note: For interpretability, the spatial distributions in the circular economy are visualized as if it is on a line segment. In each figure, the leftmost region is neighboring to the rightmost one under Assumption C.
Figure 9A considers the Type L model by Helpman-Book1998. The symmetry $\bar{\bm{x}}$ is unstable if $\phi$ is small, and agents concentrate around a single peak. As $\phi$ increases, it causes the spreading of the single agglomeration as the peak population density is reduced. At a critical level of $\phi$, the spatial distribution converges to $\bar{\bm{x}}$ [Proposition 1 (a)]. In particular, stable equilibria are single-peaked throughout the process.
Figure 9B considers Krugman-JPE1991’s model as a leading example of Type G. At low levels of $\phi$, the uniform distribution is stable. As we increase $\phi$, an endogenous transition occurs at a critical threshold of $\phi$ [Proposition 1 (b)]: a polycentric pattern with eight agglomerations becomes a stable equilibria. This eight-centered configuration remains stable over a range of $\phi$. Further increases in $\phi$ reduce the number of agglomerations, widen the spacing between them, and enlarge each center, via successive instability of equilibria similar to Proposition 3. The number of economic centers in stable spatial equilibrium evolves as $16\to 8\to 4\to 2\to 1$, reflecting fewer and more widely spaced agglomerations as transport costs fall. [^19] In this class of models, established centers may decline because the growth of larger agglomerations comes at the expense of smaller ones. For example, in Fig. 9B, the fifth region from the left initially gains population as transport improves but later loses population due to increasing concentration elsewhere.
Type LG combines the properties of Types L and G. Figure 9C considers a Type LG model proposed by Pfluger-Suedekum-JUE2008. The uniform distribution is stable when $\phi$ is small. As $\phi$ increases, octacentric agglomerations emerge at some point as in Type G models. Multiple bell-shaped economic agglomerations are generated at moderate levels of transport costs. Further increase in $\phi$ causes a decrease in the number of agglomerations (as in Type G) and the spreading of each agglomeration (as in Type L). When $\phi$ is close to one, the spatial distribution becomes monocentric, similar to Type L. At the low transport cost regime, in particular, the model exhibits a transformation from a two-peaked to a single-peaked pattern accompanied by local spreading, which resembles the dual evolution of cities, i.e., global concentration and local flattening, discussed in Section 1. While the real-world dynamics appear to unfold simultaneously, the overall evolutionary process in the model generates these changes somewhat sequentially: first through a reduction in the number of peaks, and then through the flattening of the remaining agglomeration. This discrepancy may reflect limitations of the model, the most fundamental of which would be its static structure.
4   Concluding discussions
This study illustrated how the spatial scale—local or global—of dispersion forces embedded in a many-region spatial model may lead to qualitatively different implications on equilibrium spatial patterns of economic activities and their comparative statics. Leveraging the tractability of the symmetric circular geography, this study provides a conceptual lens through which the distinction between local and global forces can be formalized, offering a general framework for interpreting results across theoretical, quantitative, and empirical research. The robustness of our theoretical conclusions beyond the circular economy is a crucial issue. Below, we provide brief explorations on this respect with discussion on related literature, leaving more thorough and rigorous analysis for future research.
[…]
A The Allen–Arkolakis model (Type L)
B The Krugman model (Type G)
Figure 10: Stable spatial patterns in a square economy ($819^{2}=$ 81 locations).
Geographical advantages. The insights from the circular geography appear to qualitatively generalize to other setups, in which the shape of network produces innate geographical advantages. For illustration, Figure 10 compares stable equilibria for models from Types L and G in a square geography for different level of transport costs, where local regional fundamentals are homogeneous. In this geography, central regions have better accessibility compared with the peripheral regions. Similarities between Fig. 10 and Figs. 9A and 9B are evident. In particular, the spatial distribution of agents become monocentric in Type L, whereas it becomes polycentric in Type G. See Appendix D for discussions. To obtain formal results in such geographical settings, a feasible strategy would be to select a few tractable but representative models for each model type and then identify general insights when the network structure systematically varies, as Matsuyama-RIE2017 did in the context of a trade model.
Local advantages. Instead of incorporating advantages that stem from the underlying network structure, one can introduce regional heterogeneity while retaining the circular network. Appendix E formally examines how small differences in intrinsic local characteristics, such as productivity or amenities, affect the spatial equilibrium and its sensitivity to interregional transport costs. Specifically, it shows that the model’s internal dispersion mechanism plays a crucial role in determining whether the effects of such local advantages are amplified or mitigated as transport accessibility improves.
[…]
The magnitude of $\rho$ at a given transport cost level captures how the model’s endogenous forces translate variation in $\bm{a}$ into variation in regional population distribution. The key insight from Appendix E (see Propositions 4 and 5) is that the direction of $\rho$’s response to changes in transport costs is largely determined by the type of dispersion force embedded in the model. In models with global dispersion mechanisms, improved interregional access tends to magnify initial local advantages (i.e., $\rho$ increases), as population concentrates further in already favored regions. By contrast, in models with local dispersion mechanisms, the same improvement tends to dampen the effects of innate heterogeneity, leading to a more even population distribution (i.e., $\rho$ decreases). These patterns are also consistent with the simple examples shown in Figs. 3D and 3C.
A Observed (from 2020 to 1970)
B Removal of highways (Type L)
C Removal of highways (Type LG)
Figure 11: Actual and counterfactual population changes of Japanese regions.
Note: The figures are reproduced from Sugimoto-etal-DP2025. Panel (A) shows the actual population changes from 2020 to 1970 (i.e., backward in time). Panel (B) applies a many-region version of Helpman-Book1998 (Type L). In Panel (C), Sugimoto-etal-DP2025 proposes a Type LG model by extending the Helpman framework by incorporating land (as in Pfluger-Tabuchi-RSUE2010) and intermediate goods (as in Krugman-Venables-EER1996) as additional inputs in production. Importantly, the Type LG model does not assume immobile workers. Instead, land as the immobile factor produces a global dispersion force, as in Pfluger-Tabuchi-RSUE2010. For further details and discussion, see Sugimoto-etal-DP2025.
Geographical and local advantages. The type of dispersion mechanism can lead to markedly different comparative statics even in real-world geographies with both geographical and local heterogeneity. Figure 11 illustrates observed and counterfactual population changes across Japanese regions, focusing on the long-run impact of highways. Panel (A) shows actual changes from 2020 to 1970 (i.e., backward in time), highlighting the historical shift toward central regions over the past 50 years. Panels (B) and (C) present counterfactual simulations using the calibrated Type L and Type LG models, respectively. Each model is calibrated to the 2020 population distribution and used to assess the effect of removing the highway network (Fig. 1A).
While neither model is expected to replicate the actual evolution of population patterns, a striking contrast emerges between the two. In the Type L model, which features only local dispersion forces, the removal of highways leads to further concentration in core regions such as Tokyo. This behavior—greater centralization as transport costs rise—is consistent with the theoretical properties of Type L models. By contrast, the Type LG model, which incorporates both local and global dispersion forces, produces the opposite outcome: the same highway removal triggers a substantial population shift toward peripheral regions. This qualitative reversal highlights how the nature of embedded dispersion forces can fundamentally shape counterfactual implications in quantitative applications.
These observations call for further research on the role of endogenous economic forces in quantitative spatial models (QSMs). Redding-Rossi-Hansberg-ARE2017 states that the literature “does not aim to provide a fundamental explanation for the agglomeration of economic activity, but rather to provide an empirically relevant quantitative model to perform general equilibrium counterfactual policy exercises” (p.23). Building on this premise, the literature interprets unexplained interregional variation in endogenous outputs of the model (i.e., structural residuals) as innate location fundamentals. Hence, “agglomeration in these models is simply the result of exogenous local characteristics augmented by endogenous economic mechanisms” (Ibid.), and endogenous forces have played less featured roles than they did in stylized theories of spatial agglomeration. [^20] In Redding-Sturm-AER2008 and Allen-Arkolakis-QJE2014, the replication data indicate that the structural residuals (interpreted as housing stock in the former, and unobserved amenity and productivity in the latter) account for 90 % and 78 % of the log-city size variation in their context. However, as demonstrated by the simple experiment in Fig. 11, endogenous dispersion mechanisms can have a significant qualitative impact on counterfactual results in QSMs and hence policy suggestions.
Baum-Snow-etal-JUE2020 notes that, in the context of Chinese urban system, “the presence of an urban hierarchy appears to require either returns to scale that are important enough to permit multiple equilibria as in Krugman-JPE1991, or else an important role for industrial specialization that depends on prefecture abundance of land or natural resources,” highlighting the possible importance of strong endogenous forces as well as immobile factors. That said, an exciting yet challenging quantitative question for future research would be whether placing greater emphasis on endogenous economic forces enhances the explanatory power and/or empirical fit of QSMs.
Generally, spatial models that emphasize only one type of dispersion force might capture only certain aspects of how economic activity responds to policy and technological changes. Notably, most regional-scale quantitative models emphasize local dispersion force as the spreading mechanism, and thus tend to predict population decentralization in response to reduced transport costs, as Fig. 11B illustrates. Our results suggest that, to capture the differential regional impacts—the winners and losers—due to new large-scale transport infrastructures, accounting for global dispersion forces can be crucial, warranting further theoretical exploration of their microfoundations. See also Graham-Horcher-RP2024 for a thoughtful evaluation of conventional QSMs as potential tools for ex-ante welfare analysis of transport interventions in practice.
Empirical connections. Notably, our framework provides a unifying and coherent interpretation of various empirical findings in the literature. For example, the evidence on regional agglomeration presented by Duranton-Turner-REStud2012 and Faber-REStud2014 aligns with the predictions of Type G models. The former focused on the growth of large metropolitan areas in the United States, while the latter examined the growth of peripheral counties in China. The former study revealed a positive correlation between regional growth rates and the level of transport infrastructure, whereas the latter found a negative correlation. These opposite responses can be interpreted as different sides of the same coin for Type G models: both results indicate a tendency of concentration towards bigger regions when interregional transport access improves. Baum-Snow-etal-JUE2020 also provided an evidence for global concentration across Chinese prefectures due to the introduction of the Chinese national highway system by showing that hinterland prefectures experienced slower growth rates compared to regional primates.
Internal structure of cities and intercity distribution of economic activities are often discussed separately, and an interesting subtlety emerges here. The number of interregional highway linkages (e.g., the number and total length) within a given region is often used as a measure of intra-urban transport infrastructure, including Duranton-Turner-REStud2012 discussed above. Here, we interpret this variable instead as an indicator of exposure to improvements in the inter-regional transport network. This interpretation allows an alternative explanation of their results in the context of Type G models. This highlights the importance of a unified analytical perspective that bridges interregional and intraurban spatial structures, where the same empirical measure may carry distinct theoretical meanings depending on the spatial scale of interpretation.
The behavior of Type L or LG models corroborates with empirical evidence for the local dispersion of economic agglomerations when interregional access improves. Baum-Snow-QJE2007 and Baum-Snow-et-al-REStat2017 presented evidence for US metro areas during 1950–1990 and Chinese prefectures during 1990–2010, respectively. These studies addressed the changes in the population or production size of the central area within the bigger region, both reporting a negative effect of improved interregional access. The local spreading of cities can also be interpreted as suburbanization in response to the improved intraurban transport infrastructure or diffusion of motorized transportation in classical Alonso–Muth–Mills theory. Structural change away from agriculture also frees up land around cities and contributes to urban density decline (Coeurdacier-etal-DP2022), which is also consistent with lowering the opportunity cost of land in the Alonso–Muth–Mills framework.
Future directions. We conclude by identifying several natural extensions of our framework that warrant further theoretical investigation.
First, it is important to consider multiple types of mobile agents that differ in their proximity matrices and/or the degree of increasing returns they experience. Such heterogeneity is ubiquitous in multi-sector models (Fujita-et-al-EER1999; Hsu-EJ2012; Gaubert-AER2018; Davis-Dingel-JIE2020) and in intracity models with multiple types of agents (e.g., Fujita-Ogawa-RSUE1982; Lucas-Rossi-Hansberg-ECTA2002; Ahlfeldt-et-al-ECTA2015; Heblich-etal-QJE2020). Duranton-Morrow-Turner-REStud2014 studied the impact of new highway connections on intercity trade in the US and showed that heavier industries are more sensitive to improved access. The circular geography provides a tractable starting point for many-location analysis under such structures (Tabuchi-Thisse-JUE2011; Osawa-Akamatsu-DP2019). As Hsu-EJ2012 suggests, incorporating sectoral heterogeneity is also important for understanding the mechanisms behind the remarkable regularities in the size and spatial variation of cities (Mori-etal-PNAS2020).
Second, models with a continuum of agents are complementary to granular spatial models (e.g., Ahlfeldt-et-al-CEPR2022), in which endogenous agglomeration arises from increasing returns and the indivisibility of agents. While continuum models can replicate systematic spatial regularities—such as periodic agglomeration patterns and city-size distributions including their fractal structure (e.g., Hsu-EJ2012; Tabuchi-Thisse-JUE2011; Mori-et-al-DP2022)—granular spatial models are better suited to capturing idiosyncratic location choices by superstar firms and large plants (e.g., Greenstone-et-al-JPE2010). Combining these two approaches may yield a deeper understanding of the spatial patterns of economic activities as outcomes of endogenous mechanisms.
Appendix A Proofs
A.1   Proof of Lemma 1
Consider an infinitesimal migration shock at the uniform distribution $\bar{\bm{x}}$ so that the spatial distribution becomes $\bar{\bm{x}}+\bm{z}$. Let $T\equiv\{\bm{z}\in\mathbb{R}^{N}\mid\sum_{i\in\mathcal{I}}z_{i}=0\}$ be the set of all feasible deviations that preserves the total population.
The aggregate utility gain for migrants under a given deviation $\bm{z}$ can be measured by $\omega(\bm{z})\equiv\bm{z}^{\top}\mathbf{V}\bm{z}$, where $\top$ denotes transpose. To see this, let $\delta v_{i}(\bm{z})\equiv v_{i}(\bar{\bm{x}}+\bm{z})-v_{i}(\bar{\bm{x}})$ be the utility difference in each region under deviation $\bm{z}$. Then, the utility gain for a migrant from region $j$ to $i$ is $\frac{\bar{x}}{\bar{v}}\left(\delta v_{i}(\bm{z})-\delta v_{j}(\bm{z})\right)$. By adding up the utility gains for all the migrants in the economy, the aggregate utility gain for migrants is measured by $\omega(\bm{z})=\frac{\bar{x}}{\bar{v}}\sum_{i\in\mathcal{I}}\delta v_{i}(\bm{z})z_{i}=\frac{\bar{x}}{\bar{v}}\delta\bm{v}(\bm{z})^{\top}\bm{z}$, as there are $z_{i}$ agents migrated to region $i$ (if $z_{i}>0$) or from region $i$ (if $z_{i}<0$), each experiencing utility difference $\delta v_{i}(\bm{z})$ at both their origin and destination. The first-order approximation $\delta\bm{v}(\bm{z})\approx\bm{v}(\bar{\bm{x}})+\nabla\bm{v}(\bar{\bm{x}})\bm{z}-\bm{v}(\bar{\bm{x}})=\nabla\bm{v}(\bar{\bm{x}})\bm{z}$ gives $\omega(\bm{z})=\bm{z}^{\top}\mathbf{V}\bm{z}$.
The aggregate utility gain determines the stability of the symmetric equilibrium $\bar{\bm{x}}$. If $\omega(\bm{z})<0$ for all possible deviations, $\bar{\bm{x}}$ is stable; this is related to the concept of “regular ESS” (Sandholm-TE2010). If $\omega(\bm{z})>0$ for some deviation $\bm{z}$, $\bar{\bm{x}}$ is unstable. Thus, the maximum possible utility gain due to small migration shock is of interest.
[…]
A.2   Proof of Proposition 1
As discussed in the main text, we consider spatial models described by a payoff function (i.e., indirect utility) $\bm{v}(\bm{x})\equiv(v_{i}(\bm{x}))_{i\in\mathcal{I}}$, parametrized by a proximity matrix $[\phi_{ij}]$, along with Assumption C. Let $\mathbf{D}$ be the row-normalized proximity matrix, whose $(i,j)$th element is $\frac{\phi_{ij}}{\sum_{k\in\mathcal{I}}\phi_{ik}}$. Throughout, we assume that the payoff function $\bm{v}$ is differentiable if $x_{i}>0$ for all $i\in\mathcal{I}$, which holds true for all examples discussed in this paper. The precise version of the symmetry of exogenous local fundamentals in Assumption C is the following:
Assumption S.
For all $\bm{x}$, payoff function $\bm{v}$ satisfies $\bm{v}(\mathbf{P}\bm{x})=\mathbf{P}\bm{v}(\bm{x})$ for all permutation matrices $\mathbf{P}$ that satisfy $\mathbf{P}\mathbf{D}=\mathbf{D}\mathbf{P}$. ∎
Example 1.
Suppose $N=4$. For instance, consider regions 1 and 3. Then, swapping their indices corresponds to applying the following permutation matrix:
$\mathbf{P}=\begin{bmatrix}0&0&1&0\\ 0&1&0&0\\ 1&0&0&0\\ 0&0&0&1\end{bmatrix}.$ (A.1)
This matrix simply switches the values of $x_{1}$ and $x_{3}$ in any vector $(x_{1},x_{2},x_{3},x_{4})$. That is, $\mathbf{P}\bm{x}=(x_{3},x_{2},x_{1},x_{4})$. This corresponds to relabeling the regions while keeping their physical positions fixed. The condition $\mathbf{P}\mathbf{D}=\mathbf{D}\mathbf{P}$ ensures that the relabeling preserves the spatial relationships encoded in $\mathbf{D}$. If $\bm{v}$ does not include any region-specific heterogeneities, then the transformed utility vector must satisfy $\bm{v}(\mathbf{P}\bm{x})=\mathbf{P}\bm{v}(\bm{x})$. This property is called equivariance; it ensures that utility differences are determined entirely by the spatial distribution, not by arbitrary index labels. Equivariance allows us to employ the machineries from group-theoretic bifurcation theory (see, e.g., Golubitsky-Stewart-Book2003; Golubitsky-et-al-Book2012; Ikeda-Murota-Book2014). ∎
Under Assumption C, we can use the full dispersion as the initial state.
Lemma 3.
[…]
Proof.
For any permutation matrix $\mathbf{P}$, $\bar{\bm{x}}=\mathbf{P}\bar{\bm{x}}$. Then, $\bm{v}(\mathbf{P}\bar{\bm{x}})=\mathbf{P}\bm{v}(\bar{\bm{x}})$ reduces to $\bm{v}(\bar{\bm{x}})=\mathbf{P}\bm{v}(\bar{\bm{x}})$ for all permutation matrix $\mathbf{P}$ that satisfies $\mathbf{P}\mathbf{D}=\mathbf{D}\mathbf{P}$. This implies that $v_{i}(\bar{\bm{x}})=v_{j}(\bar{\bm{x}})$ for any $i,j\in\mathcal{I}$. That is, $\bar{\bm{x}}$ is a spatial equilibrium. ∎
We focus on a class of models that include all models discussed in the main text. As in the main text, let $\mathbf{V}=\frac{\bar{x}}{\bar{v}}\nabla\bm{v}(\bar{\bm{x}})$ be the benefit matrix for a given payoff function.
Definition 1.
A canonical model is a model associated with a rational function $\Omega$ that is continuous over $[0,1]$ such that $\mathbf{V}=\Omega\left(\mathbf{D}\right)$. We call $\Omega$ the gain function of the model.
In Definition 1, a rational function $\Omega$ is a function of form $\Omega(\cdot)=\Omega^{\sharp}(\cdot)/\Omega^{\flat}(\cdot)$ with polynomials $\Omega^{\sharp}(\cdot)$ and $\Omega^{\flat}(\cdot)\neq 0$, where our convention is that $\Omega^{\flat}(\cdot)>0$. Given such $\Omega$, we let $\Omega(\mathbf{D})=\Omega^{\flat}(\mathbf{D})^{-1}\Omega^{\sharp}(\mathbf{D})$, where, for a polynomial $P(\Theta)=c_{0}+c_{1}\Theta+c_{2}\Theta^{2}+\cdots$, we define $P(\mathbf{D})=c_{0}\mathbf{I}+c_{1}\mathbf{D}+c_{2}\mathbf{D}^{2}+\cdots$, with $\mathbf{I}$ being the identity matrix.
Below, we study the stability of $\bar{\bm{x}}$ in canonical models. Formally, we must introduce some dynamics to define the stability of $\bar{\bm{x}}$ and study agglomeration from there. For a wide class of dynamics, however, we can focus on the analysis of the benefit matrix $\mathbf{V}$.
Lemma 4.
[…]
Proof.
See Section A.3. ∎
We can assume, for example, the replicator dynamics (Taylor-Jorker-MB1978) to define local stability of spatial equilibria (see the proof of Lemma 4 for more examples).
Thus, we only need the eigenvalues of $\mathbf{V}$. A useful fact is that, if $\{(\Theta_{k},\bm{z}_{k})\}$ are the eigenpairs (eigenvalue–eigenvector pairs) of the normalized proximity matrix $\mathbf{D}$, then the eigenpairs of $\mathbf{V}=\Omega(\mathbf{D})$ are given by $\{(\Omega(\Theta_{k}),\bm{z}_{k})\}$ (e.g., Horn-Johnson-Book2012, Section 1.1). Thus, we need the eigenpairs of $\mathbf{D}$, and we have the following lemma.
Lemma 5 (Corollary of Akamatsu-Takayama-Ikeda-JEDC2012, Lemma 4.2).
Assume Assumption C. The largest eigenvalue of $\mathbf{D}$ is $\Theta_{0}\equiv 1$, with $\bm{z}_{0}=(1,1,\ldots,1)$ being the associated eigenvector. Including $\Theta_{0}$, there are $\frac{N}{2}+1$ distinct eigenvalues. Every eigenvalue $\Theta_{k}$ ($k\neq 0$) is a strictly decreasing function of $\phi$, with $\lim_{\phi\downarrow 0}\Theta_{k}=1$ and $\lim_{\phi\uparrow 1}\Theta_{k}=0$. Let $\Theta_{\text{max}}$ denote the largest eigenvalue and $\Theta_{\text{min}}$ denote the smallest eigenvalue of $\mathbf{D}$ excluding $\Theta_{0}$, respectively. Then,
$\Theta_{\text{max}}=\Theta_{1}\equiv\frac{1-\phi}{1+\phi}\cdot\frac{1-\phi^{2}}{1-2\cos(\kappa)\phi+\phi^{2}}\cdot\frac{1+\phi^{N/2}}{1-\phi^{N/2}}$ (A.2)
$\Theta_{\text{min}}=\Theta_{N/2}\equiv\left(\frac{1-\phi}{1+\phi}\right)^{2}$ (A.3)
[…]
Since $\Theta_{k}\in(0,1)$ for all relevant $k$ and $\Omega$ is well-defined for all $[0,1]$, the eigenpairs of $\mathbf{V}=\Omega(\mathbf{D})$ are in fact given by $\{(\Omega(\Theta_{k}),\bm{z}_{k})\}$. Thus, with $\omega_{k}\equiv\Omega(\Theta_{k})$, the symmetric equilibrium $\bar{\bm{x}}$ is stable if $\omega_{k}<0$ for all $k$.
As discussed in Section 2, if agglomeration (dispersion) force of the model is too strong, $\Omega(\Theta)>0$ ($\Omega(\Theta)<0$) can happen for all $\Theta$ whereby $\bar{\bm{x}}$ is unstable (stable) for all $\phi$. As we are interested in spatial agglomeration in the course of changing $\phi$, we assume that $\bar{\bm{x}}$ can switch its stability depending on $\phi$:
Assumption E (Endogenous agglomeration occurs).
The values of the model parameters are such that $\Omega$ switches its sign at least once in $(0,1)$. ∎
Under Assumption E, we can define three prototypical classes of canonical models (see Fig. 4 in the main text for illustration).
Definition 2.
Under Assumption E, a canonical model with gain function $\Omega$ is
  1. (a)
    Type L, if there can be one and only one $\Theta^{**}\in(0,1)$ such that $\Omega(\Theta)<0$ for $\Theta\in(0,\Theta^{**})$, $\Omega(\Theta^{**})=0$, and $\Omega(\Theta)>0$ for $\Theta\in(\Theta^{**},1)$.
  2. (b)
    Type G, if there can be one and only one root $\Theta^{*}\in(0,1)$ for $\Omega$ such that $\Omega(\Theta)>0$ for $\Theta\in(0,\Theta^{*})$, $\Omega(\Theta^{*})=0$, and $\Omega(\Theta)<0$ for $\Theta\in(\Theta^{*},1)$.
  3. (c)
    Type LG, if there can be two $\Theta^{*},\Theta^{**}\in(0,1)$ such that $\Omega(\Theta^{*})=\Omega(\Theta^{**})=0$ and $\Theta^{**}<\Theta^{*}$, with $\Omega(\Theta)<0$ for $\Theta\in(0,\Theta^{**})\cup(\Theta^{*},1)$ and $\Omega(\Theta)>0$ for Θ($\Theta\in(\Theta^{**},\Theta^{*})$.
As discussed in the main text, the classification corresponds to the composition of the consequential dispersion forces in the model. We focus on the three model classes defined above and consider the destabilization of $\bar{\bm{x}}$. There can be a fourth class of models such that $\bm{x}$ is stable for medium levels of $\Theta$ but not for small or large $\Theta$. However, we are not aware of any model that falls into this category.
As we consider canonical models, there is a rational function $\Omega(\cdot)=\Omega^{\sharp}(\cdot)/\Omega^{\flat}(\cdot)$ with some polynomials $\Omega^{\sharp}$ and $\Omega^{\flat}(\cdot)>0$. That is, $\Omega^{\sharp}(\cdot)$ determines the sign of $\omega_{k}$ and thus governs the stability of $\bar{\bm{x}}$. We will focus on $\Omega^{\sharp}$ below, and let $\omega^{\sharp}\equiv\Omega^{\sharp}(\Theta_{k})$ so that $\operatorname{sgn}[\omega_{k}]=\operatorname{sgn}[\omega^{\sharp}_{k}]$. Fig. A.1 schematically shows connections between $\{\omega_{k}^{\sharp}\}$, $\Omega^{\sharp}(\Theta)$, and $\{\Theta_{k}\}$ to help understanding the following arguments.
Type L. By definition, there is $\Theta^{**}$ such that $\Omega^{\sharp}(\Theta)<0$ for all $\Theta\in(0,\Theta^{**})$, that $\Omega^{\sharp}(\Theta^{**})=0$, and that $\Omega^{\sharp}(\Theta^{**})>0$ for all $\Theta\in(\Theta^{**},1)$. Thus, $\bar{\bm{x}}$ is stable if and only if $\Theta_{k}\in(0,\Theta^{**})$, so that $\omega^{\sharp}_{k}=\Omega^{\sharp}(\Theta_{k})<0$, for all $k$, i.e., if $\Theta^{**}>\max_{k}\Theta_{k}=\Theta_{1}$. Thus, $\bar{\bm{x}}$ is stable for all $(\phi^{*},1)$ where ϕsuperscriptitalic-$\phi^{**}$ is the unique solution for $\Theta_{1}(\phi)=\Theta^{**}$. Because $\Omega^{\sharp}(\Theta)>0$ for all $\Theta\in(\Theta^{**},1)$ and $\Theta_{1}$ is strictly decreasing, $\bar{\bm{x}}$ is unstable for all $(0,\phi^{**})$.
Type G. By definition, there is $\Theta^{*}$ such that $\Omega^{\sharp}(\Theta)<0$ for all $\Theta\in(\Theta^{*},1)$, that $\Omega^{\sharp}(\Theta^{*})=0$, and that $\Omega^{\sharp}(\Theta)>0$ for all $\Theta\in(0,\Theta^{*})$. By Lemma 5, $\{\Theta_{k}(\phi)\}$ are strictly decreasing from $1$. Thus, $\bar{\bm{x}}$ is stable if and only if $\Theta_{k}\in(\Theta^{*},1)$, so that $\omega^{\sharp}_{k}=\Omega^{\sharp}(\Theta_{k})<0$, for all $k$, i.e., if $\Theta^{*}<\min_{k}\Theta_{k}=\Theta_{N/2}$. Thus, $\bar{\bm{x}}$ is stable for all $(0,\phi^{*})$ where $\phi^{*}=(1-\sqrt{\Theta^{*}})/(1+\sqrt{\Theta^{*}})$ is the unique solution for $\Theta_{N/2}=\Theta^{*}$. Because $\Omega^{\sharp}(\Theta)>0$ for all $\Theta\in(0,\Theta^{*})$ and $\Theta_{N/2}$ is strictly decreasing, $\bar{\bm{x}}$ is unstable for all $(\phi^{*},1)$ because $\omega_{N/2}^{\sharp}>0$ for the range.
Type LG. Via similar logic, we see $\bar{\bm{x}}$ is stable if $\phi\in(0,\phi_{N/2}^{*})\cup(\phi_{1}^{*},1)$.
[…]
Figure A.1: The relationships between $\Omega^{\sharp}$, $\{\Theta_{k}\}$, and $\{\omega_{k}^{\sharp}\}$.
Note: Top: Graphs of $\omega_{k}^{\sharp}=\Omega^{\sharp}(\Theta_{k})$. Bottom left: Net gain function $\Omega^{\sharp}$ for a hypothetical Type LG model with a quadratic net gain function of the form $\Omega^{\sharp}(\Theta)=c_{0}+c_{1}\Theta+c_{2}\Theta^{2}$. Bottom right: The full set of eigenvalues $\{\Theta_{k}\}$ of $\mathbf{D}$. In the shaded regions of $\phi$ or $\Theta$, $\bar{\bm{x}}$ is stable. For the $\phi$ axis, the negative $\log$ scale is used for better readability, with the transport cost level being high toward the right. We have $\max\{\Theta_{k}\}=\Theta_{\mathrm{max}}$ and $\min\{\Theta_{k}\}=\Theta_{\mathrm{min}}$ at any given level of $\phi$.
Spatial patterns. Consider a state where $\bar{\bm{x}}$ is stable. From Lemma 4, at $\phi^{*}$ in Types G or LG, a polycentric pattern with $N/2$ peaks branches from $\bar{\bm{x}}$, while at ϕsuperscriptitalic-$\phi^{**}$ in Types L or LG, a monocentric configuration branches from $\bar{\bm{x}}$. $\square$
Remark 1.
The bifurcation towards the monocentric direction ($k=1$) is a double bifurcation, where the associated eigenvalue $\omega_{1}$ has multiplicity two—that is, two linearly independent eigenvectors exist. Migration patterns at this bifurcation take the form $c^{+}\bm{z}_{1}^{+}+c^{-}\bm{z}_{1}^{-}$ for $c^{+},c^{-}\in\mathbb{R}$. Under Assumption C, only $(c^{+},c^{-})=(c,0)$ or $(c,c)$ for some $c\in\mathbb{R}$ are admissible (ikeda2012spatial). While the choice between these two depends on the details of $\bm{v}$, this does not affect the conclusion of Proposition 1, as both combinations yield monocentric configurations. ∎
Remark 2.
Although Definition 2 introduces three prototypical model classes, Type LG models sometimes span multiple classes depending on parametric restrictions (e.g., models by Pfluger-Tabuchi-RSUE2010; Kucheryavyy-etal-JIE2024). In such cases, the parameter space can be partitioned to map model behavior to the typology. Also, the definition of models often impose parametric restrictions that fix their class. In principle, however, flexible specifications would allow empirical identification of the class supported by data through parameter estimation. This would pose a major methodological challenge, but also offers a promising direction for future research. ∎
Remark 3.
Assumption C assumes that $N$ is a multiple of four to ensure $\min_{k}\{\Theta_{k}\}=\Theta_{\frac{N}{2}}$. This is inconsequential for the broad implication of Proposition 1 on spatial patterns. If $N$ is an even, $\min_{k}\{\Theta_{k}\}=\min\{\Theta_{{\frac{N}{2}}-1},\Theta_{\frac{N}{2}}\}$. If $N$ is an odd, $\min_{k}\{\Theta_{k}\}=\min\{\Theta_{\lfloor\frac{N}{2}\rfloor},\Theta_{\lfloor\frac{N}{2}\rfloor-1}\}$. Thus, $\min_{k}\{\Theta_{k}\}$ corresponds to a polycentric direction, except for the case $N=2$ or $3$ in which polycentric patterns cannot occur. ∎
Remark 4.
Beyond the local result of the proposition, ikeda2012spatial characterized the possible equilibrium configurations and bifurcations in symmetric circular economy by group-theoretic analysis. Two formal predictions are worth mentioning.
First, no symmetry-breaking bifurcations can occur after the emergence of a single-peaked spatial pattern. For Type L models, the spatial configuration remains monocentric for the whole range of $\phi$ if the full dispersion is unstable (Figs. 8A, 1B and C.3)
The other prediction is that, if $M$ same-sized agglomerations are equidistantly placed on a circle, a symmetry-breaking bifurcation may reduce their number to $K<M$, with $K$ again dividing $N$ and the agglomerations remaining equidistant.
Proposition 1 (a) implies that Type G models yield $\frac{N}{2}$ agglomerations, with further bifurcations of the form $\frac{N}{2}\to\frac{N}{4}\to\frac{N}{8}\to\cdots\to 2\to 1$ expected if $N$ is a power of two (Akamatsu-Takayama-Ikeda-JEDC2012; ikeda2012spatial; Osawa-et-al-JRS2017). For Type L, detailed analyses (e.g., Takayama-Ikeda-Thisse-RSUE2020) confirm the emergence of single-peaked or monocentric patterns (Murata-Thisse-JUE2005; Helpman-Book1998), consistent with Proposition 1 and Fig. 9. Also, AMT-2016 formally compares Forslid-Ottaviano-JoEG2003 and Helpman-Book1998. All available formal results in the literature corroborates with Proposition 1 and the numerical examples presented in Fig. 9. ∎
A.3   Proof of Lemma 4
A myopic adjustment dynamic is a system of ordinary differential equations that describes the rate of change in the spatial distribution $\bm{x}$. Denote the dynamic that adjusts $\bm{x}$ over the set of all possible spatial distributions $\mathcal{X}\equiv\left\{\bm{x}\geq\bm{0}\mathrel{}\middle|\mathrel{}\sum_{i\in\mathcal{I}}x_{i}=1\right\}$ by $\dot{\bm{x}}=\bm{f}(\bm{x})$, where $\dot{\bm{x}}$ represents the time derivative satisfying $\sum_{i\in\mathcal{I}}\dot{x}_{i}=0$ so that the total population is invariant. For example, $\bm{f}(\bm{x})=\tilde{\bm{f}}(\bm{x},\bm{v}(\bm{x}))$ where $\tilde{\bm{f}}$ maps each pair $(\bm{x},\bm{v}(\bm{x}))$ of a state and its associated payoff to a motion vector $\dot{\bm{x}}$.
We require the following conditions on $\bm{f}$: (RS) $\bm{f}(\bm{x})=\bm{0}$ if $\bm{x}$ is a spatial distribution in which all populated regions earn the same payoff level, i.e., $v_{j}(\bm{x}^{*})=v_{k}(\bm{x}^{*})$ for all $j,k\in\{i\in\mathcal{I}\mid x_{i}^{*}>0\}$, (PC) $\bm{v}(\bm{x})^{\top}\bm{f}(\bm{x})>0$ if $\bm{f}(\bm{x})\neq\bm{0}$, and (Sym) $\mathbf{P}\bm{f}(\bm{x})=\bm{f}(\mathbf{P}\bm{x})$ for all permutation matrices $\mathbf{P}$. We call dynamics that satisfy (RS), (PC), and (Sym) admissible dynamics. The conditions (RS) and (PC) are called restricted stationality and positive correlation, respectively (Sandholm-Book2010). Also, (Sym) requires that $\bm{f}$ treats all regions symmetrically. Finally, we assume that $\bm{f}$ admits a $C^{1}$ extension to an open neighborhood of $\mathcal{X}$ in $\mathbb{R}^{N}$ to use simple derivatives.
Admissible dynamics include the Brown–von Neumann–Nash dynamic (Brown-vonNeumann-Rand1950; Nash-AM1951), the Smith dynamic (Smith-TS1984), and Riemannian game dynamics (Mertikopoulos-Sandholm-JET2018). The projection dynamic (Dupuis-Nagurney-AOR1993) and the replicator dynamic (Taylor-Jorker-MB1978) are representative instances of Riemannian game dynamics that satisfy (Sym), and are often applied for regional models.
For the uniform distribution $\bar{\bm{x}}$, (RS) implies $\bm{f}(\bar{\bm{x}})=\bm{0}$, i.e., $\bar{\bm{x}}$ is a stationary point of $\bm{f}$. Denote the Jacobian matrix of $\bm{f}$ at $\bar{\bm{x}}$ by $\mathbf{F}=[\frac{\partial f_{i}}{\partial x_{j}}(\bar{\bm{x}})]$. Assume that $\mathbf{F}$ has no eigenvalues with zero real parts. Then, $\bar{\bm{x}}$ is linearly stable if all the eigenvalues of $\mathbf{F}$, which we denote by $\{\eta_{k}\}$, have negative real parts, and linearly unstable if some eigenvalue has positive real parts (see,e.g., Hirsch-et-al-Book2012). Spatial equilibrium $\bar{\bm{x}}$ is said to be stable (unstable) if it is linearly stable (unstable) under admissible dynamics. The marginal case in which the largest eigenvalue has zero real parts is unimportant as it often corresponds to measure-zero subsets of the parameter space.
Under admissible dynamics, the stability of $\bar{\bm{x}}$ can be determined by $\mathbf{V}$, i.e., without checking $\mathbf{F}$ explicitly. To exclude degenerate cases, assume that there is no other equilibrium in the neighborhood of $\bar{\bm{x}}$. Then, (PC) implies that there is a neighborhood $\mathcal{O}\subset\mathcal{X}$ of $\bar{\bm{x}}$ such that $\bm{v}(\bm{x})^{\top}\bm{f}(\bm{x})>0$ for all $\bm{x}\in\mathcal{O}\setminus\{\bar{\bm{x}}\}$. For small deviation $\bm{z}=\bm{x}-\bar{\bm{x}}$ with $\bm{x}\in\mathcal{O}\setminus\{\bar{\bm{x}}\}$, $\bm{f}(\bm{x})\approx\bm{f}(\bar{\bm{x}})+\nabla\bm{f}(\bar{\bm{x}})\bm{z}=\mathbf{F}\bm{z}$, $\bm{v}(\bm{x})\approx\bm{v}(\bar{\bm{x}})+\nabla\bm{v}(\bar{\bm{x}})\bm{z}=\bar{v}\bm{1}+\frac{\bar{v}}{\bar{x}}\mathbf{V}$, and $0=\bm{1}^{\top}\dot{\bm{x}}=\bm{1}^{\top}\bm{f}(\bm{x})\approx\bm{1}^{\top}\mathbf{F}\bm{z}$. Combined together, for all $\bm{x}\in\mathcal{O}\setminus\{\bar{\bm{x}}\}$,
$\bm{v}(\bm{x})^{\top}\bm{f}(\bm{x})\approx\left(\bm{v}(\bar{\bm{x}})+\nabla\bm{v}(\bar{\bm{x}})\bm{z}\right)^{\top}\left(\bm{f}(\bar{\bm{x}})+\mathbf{F}\bm{z}\right)=\tfrac{\bar{v}}{\bar{x}}\left(\mathbf{V}\bm{z}\right)^{\top}\left(\mathbf{F}\bm{z}\right)>0.$ (A.4)
Under Assumption C, we can choose the same set of eigenvectors for $\mathbf{V}$ and $\mathbf{F}$ because they are both symmetric circulant matrices. Let $\{\bm{z}_{k}\}$ be the set of eigenvectors and let $\omega_{k}$ and $\eta_{k}$ be the eigenvalues of $\mathbf{V}$ and $\mathbf{F}$ associated with $\bm{z}_{k}$, respectively. Then, for each eigenvector $\bm{z}_{k}$ except for $\bm{z}_{0}=\bm{1}$, Eq. A.4 yields
$\left(\mathbf{V}\bm{z}_{k}\right)^{\top}\left(\mathbf{F}\bm{z}_{k}\right)=\omega_{k}\eta_{k}>0.$ (A.5)
As $\mathbf{F}$ and $\mathbf{V}$ are both symmetric, $\eta_{k}$ and $\omega_{k}$ are both real. Thus, Eq. A.5 implies $\operatorname{sgn}[\eta_{k}]=\operatorname{sgn}[\omega_{k}]$. Therefore, $\bar{\bm{x}}$ is stable under all admissible dynamics if and only if $\omega_{k}<0$ for all $k$, excluding $k=0$ that corresponds to $\bm{z}_{0}=\bm{1}$. Likewise, $\bar{\bm{x}}$ is unstable if and only if $\omega_{k}>0$ for some $k$, again excluding $k=0$.
Suppose exactly one $\omega_{k}$ changes sign from negative to positive at $\phi_{k}^{*}$. Then, from Eq. A.5, the corresponding eigenvalue of the Jacobian matrix of any admissible dynamic at $\bar{\bm{x}}$ must also cross zero at $\phi_{k}^{*}$. Bifurcation theory shows that the system departs from $\bar{\bm{x}}$ along the direction of the associated eigenvector $\bm{z}_{k}$, as it is tangent to the “unstable manifold” at the bifurcation point (see, e.g., Hirsch-et-al-Book2012; Kuznetsov-Book2013).
Appendices B to F are provided as a separate online appendix.
Appendix B Evolution of cities
B.1   Development of high-speed transport networks in Japan
Figure B.1 reports the development of high-speed railway and highway networks, respectively, in Japan between 1970 and 2020. The development of highways and high-speed railway networks in Japan was initiated by the Tokyo Olympics held in 1964. Between 1970 and 2020, the total highway (high-speed railway) length increased from 1,119 km (515 km) to 9,050 km (3,106 km), which is a more than eightfold (sixfold) increase. As both of these networks gradually expanded in the past half century, Japan can be seen as an ideal case to study the comparative statics regarding transport costs.
A Transport networks
B Network lengths
Figure B.1: The development of high-speed network in Japan
Notes: Data source: The Digital National Land Information Download Service (Highway: https://nlftp.mlit.go.jp/ksj/gml/datalist/KsjTmplt-N06-2023.html, High-speed railway: https://nlftp.mlit.go.jp/ksj/gml/datalist/KsjTmplt-N05-2023.html).
B.2   Japanese cities and their growths
We identify cities in Japan from the Grid Square Statistics of the Population Census in 1970–2020. A city is defined as an urban agglomeration (UA), which is a set of contiguous 1 km-by-1 km grid cells whose population density is at least 1000/km2 and a total population of at least 10,000. The results are robust for alternative threshold values. Figure B.2 shows the identified 431 UAs in 2020. They occupy 6% of land area while containing 80% of the national population. Populated cells with less than 1,000 residents are shown in grey, with darker grey indicating a larger population. We use only the cells reachable by roads from four major islands (Hokkaido, Honshu, Shikoku, and Kyushu). UAs are identified for each census year during the 1970–2020 period (every five years), and Unique IDs throughout the period are assigned. For the details of the algorithm to construct UAs, see Mori-Murakami-DP2025.
Figure B.2: Urban agglomerations of Japan in 2020
From 1970 to 2020, the total population in these cities grew by $55\%$, while the national population increased by only 21%. The population of the largest city, Tokyo, has grown by 67%, an increase about the same size of the second largest city, Osaka. Figures B.3 and B.6 provide closer looks. Population growth is typically associated with concurrent areal growth (Fig. B.3A), but population density generally decreased during the 50 years period (Fig. B.3B), in particular for small to medium-sized cities, indicating local spreading of these cities. It is noted that the largest cities also experienced local spreading as evident in their spatial distributions shown in Fig. 1C as well as Fig. B.4.
Figure B.3: Population and density growths from 1970 to 2020
Note: In Panel A, the dashed horizontal line indicates simple mean of the growth ratios of the identified Japanese cities from 1970 to 2020 ($1.45$), and the dot-dashed line indicates the growth ratio of the total population of Japan ($1.21$). Likewise, the dashed horizontal line in Panel B shows the arithmetic mean of the density growths ratios. In both panels, marker color encodes the area growth ratio during the same period. For the city labels, number in parentheses shows the change in a city’s population rank (or the invariant rank) from 1970 to 2020.
A The population distribution within Osaka in 1970 and 2020
B The population distribution within Nagoya in 1970 and 2020
Figure B.4: Local dispersion of Osaka (rank $=$ 2) and Nagoya (rank $=$ 3).
Note: The warmer colors indicate larger populations. The darkest grid cells have at least 20,000 inhabitants. The other thresholds are 15,000, 10,000, 5,000, 2,000, and 1,000 inhabitants.
[…]
Figure B.5: Population and area growths from 1970 to 2020
Note: The diagonal line in each panel represents the locus of unchanged population density, as it indicates $\log(\text{density ratio})=\log(\text{population ratio})-\log(\text{area ratio})=0$. Marker color encodes the population size in 2020 for the left panel, and the population density in 2020 for the right panel. The number in parenthesis after the city names represents their rank in the respective senses, and top five cities are shown.
Figure B.6: Maximum and average employment density within a city in Japan in 1975–2014
Notes: The green and blue lines show the arithmetic means of maximum and average employment densities within a city for years 1975, 1981, 1986, 1991, 1996, 2001, 2006, 2009, and 2014. The shaded area indicates the range covering 90% of the values for individual cities. The grid-cell data of employment are obtained from the Grid Square Statistics of the Census for Establishment (1975, 1981, 1986, 1991); Establishment and Enterprise census (1996, 2001, 2006); Economic Census for Business Frame (2009 and 2014) of Japan.
Employment distribution. The tendency of local dispersion is observed in alternative indicator of agglomeration other than population. Figure B.6 shows the change in the mean values of the maximum and average employment density within a city across all cities in Japan from 1975 to 2014. Their long-run trend indicates that the geographical distribution of employment in a city has flattened over the past half century.
B.3   Cities in other countries
Figure 2 in Section 1 uses Japanese Census data. To allow a parallel comparison between different countries, we employ the the LandScan™  Global Population Database, developed by the Department of Energy’s Oak Ridge National Laboratory (ORNL), as the basic grid population data to see the evolution of cities. Cities are defined in the same manner as for Japanese cities. To asses compatibility between the LandScan data and the census we compare Census and the LandScan data for Japan. LABEL:fig:Japan(0) and LABEL:fig:Japan(2) are, respectively, based on the Census and the the LandScan data. The LandScan data is only available after 2000, and are only roughly in agreement with the precise data based on Census. It is also noted that, while the LandScan data is available every year, we employ 5-year steps to avoid noises due to its data generation procedure that incorporates various interpolations. Nonetheless, LABEL:fig:gc-ld-other-countries confirms the broad tendency of nationwide concentration and local flattening of the cities. To be consistent with our theory that assume a fixed total population, we normalize the total population in each country to unity. In the cases of France, Germany and Japan, the local flattening is apparent even without the normalization.
Appendix C Eight regions
This section considers agglomeration processes in the $N=8$ circular economy. For a selected model from each model category, we follow stationary equilibria branching from $\bar{\bm{x}}$ and then numerically check the local stability of those stationary equilibrium solutions under the replicator dynamic (Taylor-Jorker-MB1978). See Appendix F for the details of the specific formulations and the parameter settings.
Type L model. Figure C.3 considers a Type L model by Allen-Arkolakis-QJE2014 (Section F.2.4). The model incorporates a local dispersion force but no global dispersion force. The uniform equilibrium $\bar{\bm{x}}$ is stable when transport costs are low (when $\phi$ is close to $1$). If we start from $\bar{\bm{x}}$ and consider the process of a monotonic decrease in $\phi$ from $\phi\approx 1$, then a unimodal pattern emerges due to the bifurcation at ϕsuperscriptitalic-$\phi^{**}$ [Proposition 1 (a)]. This is the bifurcation in the model. When $\phi$ decreases further, the spatial pattern smoothly converges to a full concentration in a single region in the lower extreme ($\phi\approx 0$). The local dispersion force is less important than the benefits of agglomeration when interregional transportation is prohibitively costly. Mobile agents prefer concentrating on a smaller number of regions because of the agglomeration forces. As $\phi$ increases, agglomeration force due to costly transportation diminishes, and the relative rise in the local dispersion force induces a crowding-out from the populated region to the adjacent regions. As a result, the spatial pattern gradually flattens and connects to $\bar{\bm{x}}$ at ϕsuperscriptitalic-$\phi^{**}$. We can interpret the region at the mode of population distribution (region $i$ such that $x_{i}>x_{i-1}$ and $x_{i}>x_{i+1}$ where mod $N$ for indices) as the location of an agglomeration. Then, this model endogenously produces at most one agglomeration.
Type G model. Figure C.3 reports stable equilibrium patterns in the course of increasing $\phi$ for the Krugman-JPE1991 model. In Fig. 1C, the black solid (dashed) curves depict the stable (unstable) equilibrium values of $x_{i}$ at each $\phi$. Figure 1D is the schematic illustration of the stable spatial pattern on the path. The letters in Fig. 1D correspond to those in Fig. 1C. The global dispersion force in the Krugman model stems from competition between firms over consumers’ demand. If $\phi$ is low (if transport costs are high), firms have few incentives to agglomerate, and the uniform distribution is stable. If we increase $\phi$, competition with firms in other regions becomes fiercer, as the markets of other regions become closer. At some point firms are better off forming small agglomerations so that each agglomeration has its dominant market area but is relatively remote from other agglomerations of firms. At the so-called “break point” $\phi^{*}$, a bifurcation from $\bar{\bm{x}}$ occurs and the spatial pattern is pushed towards the formation of $\frac{8}{2}=4$ distinct agglomerations [Proposition 1 (b)]. A further increase in $\phi$ causes the second and third bifurcations at ϕsuperscriptitalic-$\phi^{**}$ and ϕsuperscriptitalic-$\phi^{***}$, respectively. These bifurcations sequentially double the spacing between agglomerations, each time halving their number, $4\to 2\to 1$, in a close analogy to the first bifurcation at $\phi^{*}$. We can formally analyze the successive bifurcations if we assume a specific model (ikeda2012spatial; Akamatsu-Takayama-Ikeda-JEDC2012; Osawa-et-al-JRS2017, see, e.g.,). At the higher extreme of $\phi$, agents concentrate in a single region. This behavior can be understood as a gradual extension of the market area of each agglomeration.
Type LG model. With both local and global dispersion forces, Type LG models exhibit an interplay between the number of agglomerations, spacing between them (as in Type G models), and the spatial extent of each agglomeration (as in Type L models). Figure 2A shows the evolution of the number of agglomerations in the course of increasing $\phi$ under the Pfluger-Suedekum-JUE2008’s model (Section F.2.3). The number of agglomerations in a spatial distribution is defined by that of the local maxima therein. Figure 2A exhibits the mixed characteristics of Figs. C.3 and C.3, as expected. When $\phi<\phi^{*}$ or $\phi>\phi^{**}$, $\bar{\bm{x}}$ is stable. We interpret the number of agglomerations in $\bar{\bm{x}}$ as either $8$ (for a low $\phi$) or $1$ (for a high $\phi$) to acknowledge that $\bar{\bm{x}}$ at the low and high levels of $\phi$ are distinct. When $\phi$ gradually increases from $\phi\approx 0$, the number of agglomerations reduces from $8\to 4\to 2\to 1$ as in the Type G models (Fig. C.3), whereas it is always $1$ in the latter stage as per the Type L models (Fig. C.3). The initial stage is governed by a decline in the global dispersion force, while the later stage is marked by a relative rise of the local dispersion force.
A Bifurcation diagram
B Spatial patterns
Figure C.1: Type L model (Allen-Arkolakis-QJE2014)
C Bifurcation diagram
D Spatial patterns
Figure C.2: Type G model (Krugman-JPE1991)
A Number of agglomerations
B Spatial pattern
Figure C.3: Type LG model (Pfluger-Suedekum-JUE2008)
Figure 2B illustrates the spatial patterns associated with Fig. 2A. Uniform pattern $\bar{\bm{x}}$ is initially stable (Pattern A) and the first bifurcation at $\phi^{*}$ leads to a quad-modal agglomeration (B, C), whereas the second bifurcation to the formation of a bimodal agglomeration (D, E). These transitions are in line with Fig. C.3 and are governed by the gradual decline in the global dispersion force. A further decline in the global dispersion force increases the relative importance of the local dispersion force. As a result, the bimodal agglomeration flattens out gradually (F, G). When $\phi$ increases further, it reduces to a unimodal agglomeration (J, K). The unimodal agglomeration flattens out as $\phi$ increases (L, M) until it converges to the complete dispersion (N) at ϕsuperscriptitalic-$\phi^{**}$.
Appendix D Geographical advantages
The implications of Propositions 1, 2 and 3 qualitatively generalize to different geographical setups such as one-dimensional line segment, two-dimensional spaces with and without boundaries. In particular, the spatial distribution of agents become polycentric in Type G models, whereas it becomes monocentric in Type L models.
The simplest way to introduce geographical asymmetry into our one-dimensional setup is to consider a bounded line segment, which is a standard stylized setting in urban economic theory. Ikeda-et-al-IJET2017 considered a Type G model (Forslid-Ottaviano-JoEG2003) in a line segment. They showed that multiple agglomerations emerge as in the circular economy and demonstrated that the evolution of spatial structure in a line segment approximately follows the “period doubling” behavior (Akamatsu-Takayama-Ikeda-JEDC2012; Osawa-et-al-JRS2017). For Types L and LG, Fig. D.1 reports examples of endogenous agglomeration patterns in the models by Helpman-Book1998 and Pfluger-Suedekum-JUE2008. For both models, qualitative properties of the spatial patterns are consistent with those discussed in Appendix C.
The two-dimensional counterpart of the symmetric circle is bounded lattices with periodic boundary conditions, for which a basic theory of spatial agglomeration is provided in Ikeda-Murota-Book2014. For Type G models, they typically produce multiple disjointed agglomerations and period-doubling behavior as discussed in Appendix C (see, e.g., Ikeda-etal-IJBC2012; Ikeda-etal-JEBO2014; Ikeda-et-al-JRS2017; Ikeda-etal-JEDC2018). As concrete examples, Figure 10 in the main text shows endogenous equilibrium spatial patterns over a bounded square economy with $9\times 9=81$ regions in the course of increasing $\phi$ for the Krugman and Allen–Arkolakis models. The parameters are the same as Figs. C.3 and C.3. Their agglomeration processes are qualitatively consistent with Propositions 1, 2 and 3 and examples in Appendix C, suggesting the robustness of qualitative implications of our theoretical developments.
The implications of Proposition 1 seem to extend to different assumptions on transport technology that are not formally covered by Assumption C. For example, linear transport costs are often assumed in the literature (e.g., Mossay-Picard-JET2011; Picard-Tabuchi-JET2013; Blanchet-et-al-IER2016). Mossay-Picard-JET2011 considered a variant of the Beckmann model (Type L) and showed that the only possible equilibrium is a unimodal distribution in a continuous line segment. Blanchet-et-al-IER2016 considered a general Type L model over a continuous one- or two-dimensional space; they showed that the equilibrium spatial pattern for the Beckmann model is unique and given by a regular concave paraboloid, i.e., a unimodal pattern. Picard-Tabuchi-JET2013 also considered a Type L general equilibrium model in a two-dimensional space and showed that spatial distribution becomes unimodal. The numerical results of anas1996general and anas1998urban in line segments bear a close resemblance to, respectively, agglomeration behaviors of Type L and G models, although they assume endogenous transport costs between locations.
Note: There are no asymmetries in regional characteristics except for geographical accessibility. The transport cost between every consecutive pair of regions is uniform. The level of transport cost monotonically decreases from top to bottom. Panels (A) and (B) consider a line segment with 65 locations. Panel (B) adapted from Ikeda-et-al-IJET2017 considers 17 locations; see the original paper for an extensive discussion on the behavior of a Type G model in a line segment.
[…]
A Type L
B Type G
C Type LG
Figure D.1: Stable spatial patterns in a line segment.
Appendix E Local advantages
This appendix considers small asymmetries in local regional characteristics, while retaining the circular geographical structure. In particular, we show that if interregional access improves, the regions with higher exogenous local advantage experience population increase under a global dispersion force, and converse is true under a local dispersion force.
E.1   Evaluating the impacts of local characteristics
Consider a spatial model with the indirect utility function $\bm{v}$. Let $a_{i}>0$ denote the innate characteristics of region $i$, and write $\bm{a}=(a_{i})_{i\in\mathcal{I}}$. We assume there is no other source of regional heterogeneity. For example, $a_{i}$ may represent the level of exogenous amenities or productivity in region $i$ [^21] All endogenous mechanisms related to agents’ spatial distribution $\bm{x}$, including endogenous amenities (e.g., Diamond-AER2016), are embedded in the indirect utility function $\bm{v}$ in our framework..
[…]
Example 2.
In Allen-Arkolakis-QJE2014, the indirect utility takes the form
[…]
where $\bm{w}=(w_{i}){i\in\mathcal{I}}$ is the wage vector determined by the market-clearing conditions:
$w_{i}x_{i}=\sum_{j\in\mathcal{I}}\frac{w_{i}^{1-\sigma}\big{(}\overbrace{b_{i}x_{i}^{\alpha}}^{\mathclap{\text{Local productivity spillover (agglomeration force)}}}\big{)}^{\sigma-1}\phi_{ij}}{\sum_{k\in\mathcal{I}}w_{k}^{1-\sigma}\left(b_{k}x_{k}^{\alpha}\right)^{\sigma-1}\phi_{kj}}w_{j}x_{j}\quad\forall i\in\mathcal{I},$ (E.2)
with parameters $\alpha>0$, $\beta<0$, and $\sigma>1$. Here, $u_{i}$ and $b_{i}$ are region-specific primitives interpreted as innate amenities and productivity, respectively. When $\alpha+\beta<0$, the equilibrium is unique. In that case, given an observed population distribution $\bm{x}$, the values of $u_{i}$ and $b_{i}$ that rationalize $\bm{x}$ as an equilibrium are uniquely identified from the equilibrium conditions and interpreted as the regions’ innate characteristics. ∎
When equilibrium is unique, counterfactual analysis proceeds by examining how the population distribution $\bm{x}$ responds to changes in interregional transport costs, holding the calibrated regional characteristics fixed. This assumption that $\bm{a}$ remains unchanged under the counterfactual shocks is central to quantitative predictions. Hence, understanding how $\bm{a}$ influences model outcomes is critical.
Suppose now that $\bm{a}$ deviates slightly from $\bar{\bm{a}}$, so that $\bar{\bm{x}}$ is no longer an equilibrium. If the deviation is small, the resulting equilibrium $\bm{x}(\bm{a})$ will remain close to $\bar{\bm{x}}$ except for knife-edge cases. Thus, we can view $\bm{x}(\bm{a})$ as a continuous function of $\bm{a}$ satisfying $\bm{x}(\bar{\bm{a}})=\bar{\bm{x}}$.
To quantify the overall impact of variations in $\bm{a}$, we define the covariance between each region’s relative advantage and its deviation in population share from $\bar{\bm{x}}$:
$\rho\equiv\sum_{i\in\mathcal{I}}\overbrace{(a_{i}-\bar{a})}^{\mathclap{\text{Regional (dis)advantage}}}\underbrace{(x_{i}(\bm{a})-\bar{x})}_{\mathclap{\text{Population deviation}}}.$ (E.3)
For example, if $\rho=0$, variations in $\bm{a}$ have no impact on the spatial distribution. We assume $\rho>0$, reflecting the natural intuition that more advantaged regions attract more population.
Importantly, the response of $\rho$ to changes in transport costs captures how the internal structure of a model shapes its counterfactual behavior. In particular, it reveals the model’s intrinsic directional bias: whether it systematically favors the concentration of population in relatively advantaged regions ($a_{i}>\bar{a}$) under transport cost shocks. An increase in $\rho$ implies that such regions gain population in the new equilibrium; a decrease implies the opposite.
E.2   Formal characterizations
We can analytically characterize the response of $\rho$ under symmetric transport cost structures.
[…]
$\rho=(a_{1}-\bar{a})(x_{1}-\bar{x})+(a_{2}-\bar{a})(x_{2}-\bar{x})=\epsilon\xi+(-\epsilon)(-\xi)=2\epsilon\xi>0.$ (E.4)
Notably, we can obtain the analytical expression for $\rho$ for a given spatial model. First, we recall that the utility gain due to migration from $\bar{\bm{x}}$ is negative ($\omega<0$) as we assume $\bar{\bm{x}}$ is stable. Next, the utility gain induced by small regional asymmetry $(\bar{a}+\epsilon,\bar{a}-\epsilon)$ can be evaluated by the following elasticity, analogous to $\omega$:
$\omega^{\natural}\equiv\frac{\bar{a}}{\bar{v}}\left(\frac{\partial{v_{1}(\bar{\bm{x}},\bar{\bm{a}})}}{\partial{a_{1}}}-\frac{\partial{v_{2}(\bar{\bm{x}},\bar{\bm{a}})}}{\partial{a_{1}}}\right),$ (E.5)
where $\bar{v}\equiv v_{i}(\bar{\bm{x}},\bar{\bm{a}})$. The dependence of $\bm{v}$ on $\bm{a}$ is made explicit. Then, we have:
Lemma 6.
Assume $N=2$. For the perturbed equilibrium under $\bm{a}=(\bar{a}+\epsilon,\bar{a}-\epsilon)$ with small $\epsilon$,
$\rho=-c\cdot\frac{\omega^{\natural}}{\omega}\quad\text{where}\quad c\equiv 2\epsilon^{2}\frac{\bar{x}}{\bar{a}}.$ (E.6)
Proof.
Let $f(x,a)\equiv v_{1}(x,a)-v_{2}(x,a)$ with $x\coloneq x_{1}$ and $a\coloneq a_{1}$. Then, $f(x,a)=0$ since the new spatial distribution $\bm{x}$ is an equilibrium. At $(x,a)=(\bar{x},\bar{a})$ and on the equilibrium curve $f(x,a)=0$, we have $0=f_{x}(x,a)\xi+f_{a}(x,a)\epsilon=\frac{\bar{x}}{\bar{v}}f_{x}(x,a)\frac{\bar{v}}{\bar{x}}\xi+\frac{\bar{a}}{\bar{v}}f_{a}(x,a)\frac{\bar{v}}{\bar{x}}\epsilon=\omega\frac{\bar{v}}{\bar{x}}\xi+\omega^{\natural}\frac{\bar{v}}{\bar{a}}\epsilon$. That is, $\omega<0$ and $\omega^{\natural}>0$ should counterweight each other. From $\rho=2\epsilon\xi$, we obtain the desired formula. ∎
Lemma 6 implies the following result on the response of $\rho$ to the increase in $\phi$.
Proposition 4.
Assume $N=2$. If local regional characteristic is multiplicatively separable as $v_{i}(\bm{x},\bm{a})=a_{i}v_{i}(\bm{x})$ with $v_{i}(\bm{x})$ satisfying Assumption S (i.e., no other exogenous asymmetries),
$\operatorname{sgn}\rho^{\prime}(\phi)=\operatorname{sgn}\omega^{\prime}(\phi).$ (E.7)
Proof.
Under the hypothesis of the claim, $\omega^{\natural}=\frac{\bar{a}}{\bar{v}}(v_{1}(\bar{\bm{x}})-0)=\bar{a}>0$. Then, Eq. E.6 reduces to $\rho^{\prime}(\phi)=\frac{c\bar{a}}{\omega^{2}}\omega^{\prime}(\phi)$ where denotes the differentiation by $\phi$, implying Eq. E.7. ∎
A The Redding–Sturm model (Type L)
B The Krugman model (Type G)
Figure E.1: The curve of the utility gain $\omega$ for the two-region case.
While the sign of Eq. E.6 is model-dependent, there is a broad connection between the spatial scale of dispersion forces and the sign of $\rho^{\prime}(\phi)$. In Type G models, agglomeration occurs when $\phi$ increases. Reflecting this, for Type G models in the literature, we have $\omega^{\prime}(\phi)>0$ if $\bar{\bm{x}}$ is stable. Converse is true for Type L models, for which dispersion occurs when transport access improves and $\omega^{\prime}(\phi)<0$ if $\bar{\bm{x}}$ is stable. Fig. E.1 shows the curves of $\omega$ for the Krugman model (Type G) and the Redding–Sturm model (Type L), which demonstrates that Types L and G can come to the opposite conclusions in the two-region economy. If a model has only a global dispersion force, its mechanisms strengthen the effects of exogenous local advantages in innate amenities, and the converse is true for a model with only a global dispersion force.
For other forms of exogenous fundamentals, such clean characterization is not available. For example, heterogeneities in innate regional productivity can affect the utility level of other regions through interregional trade, and thus are not multiplicatively separable.
The circular geography. Nonetheless, we have a characterization of the response of $\rho$ for general local characteristics. Now suppose the symmetric circle (Assumption C). Several notations and assumptions are in order. As we have seen in Section 3, the utility elasticity matrix $\mathbf{V}=\frac{\bar{x}}{\bar{v}}[\frac{\partial v_{i}}{x_{j}}(\bar{\bm{x}})]$ at $\bar{\bm{x}}$ is simply represented by the row-normalized proximity matrix $\mathbf{D}$. Suppose $\mathbf{V}=\Omega(\mathbf{D})$ where $\Omega$ is a scalar-valued rational function that is continuous over $[0,1]$ and the interpretation of $\Omega(\mathbf{D})$ is the same as in Section A.2. In an analogous manner, let $\mathbf{A}=\frac{\bar{a}}{\bar{v}}[\frac{\partial v_{i}}{\partial a_{j}}(\bar{\bm{x}})]$ be the utility elasticity matrix with respect to the local characteristic vector $\bm{a}$ of interest, and suppose that $\mathbf{A}$ can be represented by $\mathbf{D}$. Specifically, let $\mathbf{A}=\Omega^{\natural}(\mathbf{D})$ where $\Omega^{\natural}(\cdot)$ is another rational function that is continuous over $[0,1]$. In fact, regional heterogeneities in $\bm{a}$ can be seen as deviations from $\bar{\bm{a}}$, so the effects of such deviations on the utility vector can be evaluated exactly the same manner as population deviations from $\bar{\bm{x}}$ considered in the two-region case. Let
$\delta(\Theta)=-\frac{\Omega^{\natural}(\Theta)}{\Omega(\Theta)},$ (E.8)
which corresponds to $-\frac{\omega^{\natural}}{\omega}$ in Eq. E.6 for the two-region case.
Proposition 5.
Suppose Assumption C and assume that $\bar{\bm{x}}$ is stable. Then,
  1. (a)
    $\rho^{\prime}(\phi)>0$, if $\delta^{\prime}(\Theta)<0$ for all $\Theta\in(0,1)$ such that $\Omega(\Theta)<0$.
  2. (b)
    $\rho^{\prime}(\phi)<0$, if $\delta^{\prime}(\Theta)>0$ for all $\Theta\in(0,1)$ such that $\Omega(\Theta)<0$.
Proof.
See Section E.4. ∎
The multiplicatively separable case is the special case in which $\Omega^{\natural}(\cdot)=\bar{a}$. For the general case with non-constant $\Omega^{\natural}$, our three model classes are not precisely mapped to Proposition 5 (a) or (b). Still, there is a broad tendency that initial advantages are amplified in Type G models, whereas they are diminished in Type L models. For example, the regional-scale model considered in Redding-Rossi-Hansberg-ARE2017 is Type L. If we consider local productivity parameters as regional characteristic vector $\bm{a}$, then the model satisfies $\delta^{\prime}(\Theta)>0$ for all $\Theta\in(0,1)$ when equilibrium is unique (see Remark 11 in Appendix F). Likewise, the Krugman model is Type G and we show $\delta^{\prime}(\Theta)<0$ for all $\Theta\in(0,1)$ if we consider immobile demand $l_{i}$ as regional characteristics (see Remark 7 in Appendix F).
E.3   Numerical examples
This section provides numerical examples for Proposition 5. To introduce exogenous asymmetry, we multiply the utility in region $1$ by $a_{1}\geq 1$, whereas we let $a_{i}=1$ for all $i\neq 1$. We consider the Krugman model and Allen–Arkolakis model, and basic model parameters are set to be the same as Fig. C.3 and Fig. C.3 except that region $1$ has an exogenous advantage.
Figure 2A reports equilibrium paths of $x_{1}$ for the Allen–Arkolakis model (Type L) under the uniqueness of the equilibrium. The curves depict region $1$’s population share, $x_{1}$, at stable equilibria against $\phi$. Four incremental settings $a_{1}\in\{1.000,1.001,1.005,1.010\}$ are considered, including the baseline case with no location-fixed advantage ($a_{1}=1.000$). We have $\delta^{\prime}(\Theta)>0$ for all $\Theta\in(0,1)$ and see that $x_{1}-\bar{x}>0$ when $a_{1}>1$ and $x_{1}-\bar{x}$ increases as $a_{1}$ increases, which are intuitive. Additionally, $x_{1}-\bar{x}$ decreases as $\phi$ increases. We confirm that $\rho(\phi)>0$ and $\rho^{\prime}(\phi)<0$ for all $\phi$.
Figures 2C and 2B consider the Krugman and Allen–Arkolakis models under a multiplicity of equilibria, respectively. Unlike the Allen–Arkolakis model, the Krugman model admits multiple equilibria for some $\phi$ for any pair of the structural parameters $(\mu,\sigma)$. Proposition 5 correctly predicts the sign of $\rho^{\prime}(\phi)$ for the range of $\phi$ such that $\bar{\bm{x}}$ is stable when $a_{1}=1$; we have $\rho^{\prime}(\phi)>0$ when $\phi\in(0,\phi^{*})$ for the Krugman model, whereas $\rho^{\prime}(\phi)<0$ when $\phi\in(\phi^{**},1)$ for the Allen–Arkolakis model.
A Type L (Uniqueness)
B Type L (Multiplicity)
C Type G
Figure E.2: Population share of the advantageous region $1$ and covariance $\rho$
In Fig. 2C, the definition of $\rho$ is modified for spatial patterns with unpopulated regions. For the range $\phi\in(\phi^{*},\phi^{**})$, $\rho$ is evaluated with respect to the four-centric pattern $(2\bar{x},0,2\bar{x},0,2\bar{x},0,2\bar{x},0)$: $\rho\equiv\sum_{i\in\mathcal{I}(\bm{x})}(x_{i}-2\bar{x})(a_{i}-\bar{a}(\bm{x}))$, where $\mathcal{I}(\bm{x})=\{i\in\mathcal{I}\mid x_{i}>0\}$ is the set of populated regions and $\bar{a}(\bm{x})\equiv\frac{1}{|\mathcal{I}(\bm{x})|}\sum_{i\in\mathcal{I}(\bm{x})}a_{i}$. We define $\rho$ for two-centric pattern $(4\bar{x},0,0,0,4\bar{x},0,0,0)$ similarly. For the transitional phase after ϕsuperscriptitalic-$\phi^{**}$ we let
$\rho\equiv\sum_{i\in\mathcal{I}(\bm{x})}(x_{i}-x_{i}^{*})(a_{i}-\bar{a}(\bm{x})),$ (E.9)
where $x_{i}^{*}$ corresponds to the stable solution for the symmetric case ($a_{1}=1$).
For Fig. 2B, we employ Eq. E.9 as the definition of $\rho$ for the case $\phi\in(0,\phi^{**})$, i.e., we consider the deviation from the baseline equilibrium ($a_{1}=1$). We observe that $\rho^{\prime}(\phi)<0$ does not necessarily hold true for $\phi\in(0,\phi^{**})$. For instance, $\rho^{\prime}(\phi)>0$ when $\phi$ is small. Nonetheless, for the range of $\phi$ under which $\bar{\bm{x}}$ is stable, $\rho$ decreases in $\phi$, consistent with Proposition 5.
E.4   Proof of Proposition 5
We first derive the analytical expression of $\rho$. The equilibrium condition when all regions are populated is given by
$\bm{v}(\bm{x},\bm{a})-\bar{v}(\bm{x},\bm{a})\bm{1}=\bm{0},$ (E.10)
where we make the dependence of $\bm{v}$ on $\bm{a}$ explicit, $\bar{v}(\bm{x},\bm{a})\equiv\sum_{i\in\mathcal{I}}v_{i}(\bm{x},\bm{a})x_{i}$ is the average utility, and $\bm{1}$ is $N$-dimensional all-one vector. The pair $(\bar{\bm{x}},\bar{\bm{a}})$ is a solution to Eq. E.10. Suppose that there is a spatial equilibrium nearby $\bar{\bm{x}}$ when $\bm{a}$ is marginally different from $\bar{\bm{a}}$. Let $\bm{x}(\bm{a})$ denote the perturbed version of $\bar{\bm{x}}$, which is a function in $\bm{a}$. We assume $\bar{\bm{x}}$ is stable so that studying a perturbed version of it makes sense.
The covariance $\rho$ is represented as follows:
$\rho\equiv(\bm{a}-\bar{\bm{a}})^{\top}\left(\bm{x}(\bm{a})-\bar{\bm{x}}\right)=(\mathbf{C}\bm{a})^{\top}\mathbf{C}\bm{x}(\bm{a})=\bm{a}^{\top}\mathbf{C}\bm{x}(\bm{a})$ (E.11)
where $\mathbf{C}\equiv\mathbf{I}-\frac{1}{N}\bm{1}\bm{1}^{\top}$ is the centering matrix. Let $\mathbf{X}\equiv[\frac{\partial x_{i}}{\partial a_{j}}(\bar{\bm{a}})]$ be the Jacobian matrix of $\bm{x}$ with respect to $\bm{a}$ at $(\bar{\bm{x}},\bar{\bm{a}})$. Then, $\bm{x}(\bm{a})\approx\bar{\bm{x}}+\mathbf{X}(\bm{a}-\bar{\bm{a}})=\bar{\bm{x}}+\mathbf{X}\mathbf{C}\bm{a}$ and thus $\rho=\bm{a}^{\top}\mathbf{C}\mathbf{X}\mathbf{C}\bm{a}$ as $\mathbf{C}\bar{\bm{x}}=\bm{0}$. The implicit function theorem regarding Eq. E.10 at $(\bar{\bm{x}},\bar{\bm{a}})$ gives:
$\mathbf{X}$ $=-\left(\mathbf{V}_{x}-\bm{1}\bar{\bm{x}}^{\top}\mathbf{V}_{x}-\bm{1}\bm{v}(\bar{\bm{x}})^{\top}\right)^{-1}\left(\mathbf{V}_{a}-\bm{1}\bar{\bm{x}}^{\top}\mathbf{V}_{a}\right)$ (E.12)
$=\left(\tfrac{\bar{v}}{\bar{x}}\tfrac{1}{N}\bm{1}\bm{1}^{\top}-\left(\mathbf{I}-\tfrac{1}{N}\bm{1}\bm{1}^{\top}\right)\mathbf{V}_{x}\right)^{-1}\left(\mathbf{I}-\tfrac{1}{N}\bm{1}\bm{1}^{\top}\right)\mathbf{V}_{a}$ (E.13)
$=\tfrac{\bar{x}}{\bar{v}}\left((\mathbf{I}-\mathbf{C})-\mathbf{C}\tfrac{\bar{x}}{\bar{v}}\mathbf{V}_{x}\right)^{-1}\mathbf{C}\tfrac{\bar{v}}{\bar{a}}\tfrac{\bar{a}}{\bar{v}}\mathbf{V}_{a}$ (E.14)
$=\tfrac{\bar{x}}{\bar{a}}\left((\mathbf{I}-\mathbf{C})-\mathbf{C}\mathbf{V}\right)^{-1}\mathbf{C}\mathbf{A}$ (E.15)
where $\bar{v}$ is the utility level, $\mathbf{V}_{x}\equiv[\frac{\partial v_{i}}{\partial x_{j}}(\bar{\bm{x}},\bar{\bm{a}})]$, $\mathbf{V}_{a}\equiv[\frac{\partial v_{i}}{\partial a_{j}}(\bar{\bm{x}},\bar{\bm{a}})]$, $\mathbf{V}\equiv\frac{\bar{x}}{\bar{v}}\mathbf{V}_{x}$, and $\mathbf{A}\equiv\frac{\bar{a}}{\bar{v}}\mathbf{V}_{a}$.
For tractability, we focus on a specific form of $\mathbf{A}$ which covers many relevant cases.
Assumption A.
Suppose Assumption C. Let $\mathbf{A}\equiv\frac{\bar{a}}{\bar{v}}[\frac{\partial v_{i}}{\partial a_{j}}]$ be the elasticity matrix of the utility with respect to the local characteristic $\bm{a}$ under consideration, evaluated at $(\bar{\bm{x}},\bar{\bm{a}})$. There is a rational function $\Omega^{\natural}$ that is continuous over $[0,1]$, positive whenever $\bar{\bm{x}}$ is stable, and satisfies $\mathbf{A}=\Omega^{\natural}(\mathbf{D})$. ∎
Example 3.
Suppose $v_{i}(\bm{x},\bm{a})=a_{i}v_{i}(\bm{x})$, where $a_{i}>0$ is the exogenous level of local amenities and $\bm{v}(\bm{x})=(v_{i}(\bm{x}))_{i\in\mathcal{I}}$ is the symmetric component of the utility function (i.e., $\bm{v}(\bm{x})$ satisfies Assumption S). Then, $\mathbf{A}=\bar{a}\mathbf{I}$ and $\Omega^{\natural}(\cdot)=\bar{a}>0$. ∎
Under Assumptions C and A, $\mathbf{X}$ is real, symmetric, and circulant. Thus, the set of eigenvectors of $\mathbf{C}\mathbf{X}\mathbf{C}$ can be chosen as in Lemma 5 (a) because it is a circulant matrix of the same size as $\mathbf{D}$. Let $\{\lambda_{k}\}_{k=0}^{M}$ be the distinct eigenvalues of $\mathbf{C}\mathbf{X}\mathbf{C}$. As $\mathbf{C}\mathbf{X}\mathbf{C}$ is symmetric, it admits the eigenvalue decomposition
$\mathbf{C}\mathbf{X}\mathbf{C}=\lambda_{0}\bm{1}\bm{1}^{\top}+\sum_{k=1}^{M-1}\lambda_{k}\left(\bm{z}_{k}^{+}{\bm{z}_{k}^{+}}^{\top}+\bm{z}_{k}^{-}{\bm{z}_{k}^{-}}^{\top}\right)+\lambda_{M}\bm{z}_{M}\bm{z}_{M}^{\top}.$ (E.16)
This fact yields the following representation of $\rho$:
$\rho=\bm{a}^{\top}\mathbf{C}\mathbf{X}\mathbf{C}\bm{a}=\sum_{k\neq 0}\tilde{a}_{k}^{2}\lambda_{k},$ (E.17)
where $\tilde{\bm{a}}\equiv(\tilde{a}_{k})$ is the representation of $\bm{a}$ in the new coordinate system $\{\bm{z}_{k}\}$. We can drop $k=0$ because $\lambda_{0}=0$, reflecting that $\bm{z}_{0}=\bm{1}$ represents a uniform increase in $\bm{a}$ and thus does not affect spatial equilibria. All the matrices in Eq. E.15 are circulant and hence shares the same set of eigenvectors. Thus, $\lambda_{k}$ is obtained from Eq. E.15 as follows:
$\lambda_{k}=\frac{\bar{x}}{\bar{a}}\frac{\kappa_{k}\omega^{\natural}_{k}}{((1-\kappa_{k})-\kappa_{k}\omega_{k})}=-\frac{\bar{x}}{\bar{a}}\frac{\omega^{\natural}_{k}}{\omega_{k}}$ $\forall k\in\mathcal{K},$ (E.18)
where $\kappa_{k}$, $\omega_{k}$, and $\omega^{\natural}_{k}$ are the $k$th eigenvalues of $\mathbf{C}$, $\mathbf{V}$, and $\mathbf{A}$, respectively, with $\kappa_{0}=0$ and $\kappa_{k}=1$ for all $k\neq 0$. As $\omega_{k}=\Omega(\Theta_{k})$ and $\omega^{\natural}_{k}=\Omega^{\natural}(\Theta_{k})$ with $\{\Theta_{k}\}_{k\in\mathcal{K}}$ are the eigenvalues of $\mathbf{D}$ because we assume $\mathbf{G}=\Omega(\mathbf{D})$ and $\mathbf{A}=\Omega^{\natural}(\mathbf{D})$,
$\lambda_{k}=-\frac{\bar{x}}{\bar{a}}\dfrac{\Omega^{\natural}(\Theta_{k})}{\Omega(\Theta_{k})}=\frac{\bar{x}}{\bar{a}}\delta(\Theta_{k})$ $\forall k\in\mathcal{K}$ (E.19)
and $\lambda_{0}=0$ where $\delta(\Theta)\equiv-\frac{\Omega^{\natural}(\Theta)}{\Omega(\Theta)}$.
From Eq. E.17, $\rho>0$ for all $\bm{a}$ if all $\{\lambda_{k}\}$ are positive except for $\lambda_{0}=0$. The denominator of Eq. E.19, $\Omega(\Theta_{k})$, must be negative for all $k$ because $\bar{\bm{x}}$ is stable by assumption. Thus, we see that $\rho>0$ if $\Omega^{\natural}(\Theta)>0$ for all $\Theta$ since $\Theta_{k}\in(0,1)$ for all $k\in\mathcal{K}$.
Proposition 5 follows by noting
$\rho^{\prime}(\phi)=\sum_{k\neq 0}\tilde{a}_{k}^{2}\frac{\mathrm{d}{\lambda_{k}}}{\mathrm{d}\phi}=\frac{\bar{x}}{\bar{a}}\sum_{k\neq 0}\tilde{a}_{k}^{2}\delta^{\prime}(\Theta_{k})\frac{\mathrm{d}{\Theta_{k}}}{\mathrm{d}\phi}=-\frac{\bar{x}}{\bar{a}}\sum_{k\neq 0}\tilde{a}_{k}^{2}\delta^{\prime}(\Theta_{k})\left|\frac{\mathrm{d}{\Theta_{k}}}{\mathrm{d}\phi}\right|.$
From Lemma 5, $\{\Theta_{k}\}_{k\in\mathcal{K}}$ are strictly decreasing in $\phi$. Thus, for $\rho^{\prime}(\phi)>0$ ($\rho^{\prime}(\phi)<0$), it is sufficient that $\delta^{\prime}(\Theta)<0$ ($\delta^{\prime}(\Theta)>0$) for all $\Theta$ such that $\Omega(\Theta)<0$.
Appendix F Derivations
This section collects omitted derivations. Below, the expression $\mathbf{F}_{x}$ denotes the Jacobian matrix of a vector-valued function $\bm{f}(\bm{x})$ with respect to $\bm{x}$, that is, $\mathbf{F}_{x}=[\frac{\partial f_{i}}{\partial x_{j}}]$. For example, $\mathbf{V}_{x}\equiv[\frac{\partial v_{i}}{\partial x_{j}}]$, $\tilde{\mathbf{V}}_{w}\equiv[\frac{\partial\tilde{v}_{i}}{\partial w_{j}}]$, and $\mathbf{W}_{x}\equiv[\frac{\partial w_{i}}{\partial x_{j}}]$. Throughout, $\bar{v}$, $\bar{w}$, $\bar{e}$ and so on represent $v_{i}$, $w_{i}$, $e_{i}$ evaluated at $\bar{\bm{x}}$. $\mathbf{D}$ denotes the row-normalized proximity matrix.
F.1   General derivations
F.1.1   The benefit matrix
In spatial models, the payoff function (i.e., indirect utility function) $\bm{v}$ often reduce to the following implicit form:
$\bm{v}(\bm{x})=\tilde{\bm{v}}(\bm{x},\bm{w}),$ (F.1)
$\bm{s}(\bm{x},\bm{w})=\bm{0}.$ (F.2)
The condition Eq. F.2 represents, e.g., the general equilibrium conditions for a given $\bm{x}$ that defines endogenous variable $\bm{w}$ (e.g., wages) other than $\bm{x}$ as an implicit function of $\bm{x}$. For $\bm{v}(\bm{x})$ to be well-defined, Eq. F.2 must admit a unique solution of $\bm{w}$ for all $\bm{x}\in\mathcal{X}$. Suppose $\bm{s}$ and $\bm{v}$ are continuously differentiable. Then, we have
$\mathbf{V}_{x}(\bm{x})=\tilde{\mathbf{V}}_{x}(\bm{x})+\tilde{\mathbf{V}}_{w}(\bm{x})\mathbf{W}_{x}(\bm{x}),$ (F.3)
$\mathbf{W}_{x}(\bm{x})=-\mathbf{S}_{w}(\bm{x})^{-1}\mathbf{S}_{x}(\bm{x}),$ (F.4)
where $\mathbf{W}_{x}(\bm{x})$ is obtained by applying the implicit function theorem to Eq. F.2.
Under Assumption C, all relevant matrices commute at $\bm{x}=\bar{\bm{x}}$ because they are real, symmetric, and circulant at $\bar{\bm{x}}$. Thus, $\mathbf{V}_{x}=\mathbf{S}_{w}^{-1}(\mathbf{S}_{w}\tilde{\mathbf{V}}_{x}-\tilde{\mathbf{V}}_{w}\mathbf{S}_{x})$ at $\bar{\bm{x}}$.
Example 4.
Eq. F.2 is often given by
$s_{i}(\bm{x},\bm{w})=w_{i}x_{i}-\sum_{j\in\mathcal{I}}m_{ij}e_{j}=0,$ (F.5)
where regional expenditure is $e_{i}=e(w_{i},x_{i})$ with some nonnegative function $e$ and $\mathbf{M}=[m_{ij}]$ is the expenditure share matrix whose typical element being
$m_{ij}=\dfrac{x_{i}w_{i}^{1-\sigma}\phi_{ij}}{\sum_{k\in\mathcal{I}}x_{k}w_{k}^{1-\sigma}\phi_{kj}}.$ (F.6)
In matrix form, we can write $\bm{y}-\mathbf{M}\bm{e}=\bm{0}$ where $\bm{y}=(w_{i}x_{i})_{i\in\mathcal{I}}$. Then, in general,
$\mathbf{S}_{x}(\bm{x})=\operatornamewithlimits{diag}[\bm{w}]-\left(\operatornamewithlimits{diag}[\mathbf{M}\bm{e}]-\mathbf{M}\operatornamewithlimits{diag}[\bm{e}]\mathbf{M}^{\top}\right)\operatornamewithlimits{diag}[\bm{x}]^{-1}-\mathbf{M}\mathbf{E}_{x},$ (F.7a)
$\mathbf{S}_{w}(\bm{x})=\operatornamewithlimits{diag}[\bm{x}]+(\sigma-1)\left(\operatornamewithlimits{diag}[\mathbf{M}\bm{e}]-\mathbf{M}\operatornamewithlimits{diag}[\bm{e}]\mathbf{M}^{\top}\right)\operatornamewithlimits{diag}[\bm{w}]^{-1}-\mathbf{M}\mathbf{E}_{w}.$ (F.7b)
Suppose Assumption C. Suppose $\bm{x}=\bar{\bm{x}}$ and let $\bar{w}$ be the uniform level of $\{w_{i}\}$ at $\bar{\bm{x}}$. Then, we have $\mathbf{M}=\mathbf{D}$ at $\bm{x}=\bar{\bm{x}}$. Suppose the scalars $\epsilon_{x}$ and $\epsilon_{w}$ are chosen to satisfy $\mathbf{E}_{x}=\epsilon_{x}\bar{w}\mathbf{I}$ and $\mathbf{E}_{w}=\epsilon_{w}\bar{x}\mathbf{I}$ at $\bar{\bm{x}}$. Let $\bar{e}=e(\bar{w},\bar{x})$ and $\zeta\equiv\frac{\bar{e}}{\bar{w}\bar{x}}$. We see that
$\mathbf{S}_{x}=-\bar{w}\left((\zeta-1)\mathbf{I}+\epsilon_{x}\mathbf{D}-\zeta\mathbf{D}^{2}\right),$ (F.8a)
$\mathbf{S}_{w}=\bar{x}\left(\left(1+\zeta(\sigma-1)\right)\mathbf{I}-\epsilon_{w}\mathbf{D}-\zeta(\sigma-1)\mathbf{D}^{2}\right).$ (F.8b)
If $e(w_{i},x_{i})=w_{i}x_{i}$, then $\epsilon_{x}=\epsilon_{w}=1$ and $\zeta=1$, thereby $\mathbf{W}_{x}=\frac{\bar{w}}{\bar{x}}(\sigma\mathbf{I}+(\sigma-1)\mathbf{D})^{-1}\mathbf{D}$. ∎
F.1.2   The payoff elasticity matrix with respect to local characteristics
In Eq. E.15, $\mathbf{X}=[\frac{\partial x_{i}(\bar{\bm{a}})}{\partial a_{i}}]=\mathbf{X}_{a}$ acts as $\hat{\mathbf{X}}\equiv-\mathbf{V}_{x}^{-1}\mathbf{V}_{a}$ for $\bm{z}$ such that $\bm{z}^{\top}\bm{1}=\bm{0}$. Thus, $\mathbf{V}_{a}$ is of interest.
For purely local characteristics (Example 3), since $v_{i}(\bm{x},\bm{a})=a_{i}v_{i}(\bm{x})$, it follows that $\mathbf{V}_{a}=\operatornamewithlimits{diag}[\bm{v}(\bm{x})]$. At $\bar{\bm{x}}$, we have $\mathbf{V}_{a}=\bar{v}\mathbf{I}$. Thus, $\hat{\mathbf{X}}=-\bar{v}\mathbf{V}_{x}^{-1}$.
For regional characteristics that affect trade flows, the payoff function and the market equilibrium condition are, respectively, modified to $\bm{v}(\bm{x},\bm{a})=\tilde{\bm{v}}(\bm{x},\bm{w},\bm{a})$ and $\bm{s}(\bm{x},\bm{w},\bm{a})=\bm{0}$. By applying the implicit function theorem, we see $\mathbf{V}_{a}=\tilde{\mathbf{V}}_{a}+\tilde{\mathbf{V}}_{w}\mathbf{W}_{a}=\tilde{\mathbf{V}}_{a}-\tilde{\mathbf{V}}_{w}\mathbf{S}_{w}^{-1}\mathbf{S}_{a}$. As all matrices commute at $\bar{\bm{x}}$ under Assumption C, it is equivalent to consider
$\hat{\mathbf{X}}$ $=-\left(\tilde{\mathbf{V}}_{x}-\tilde{\mathbf{V}}_{w}\mathbf{S}_{w}^{-1}\mathbf{S}_{x}\right)^{-1}\left(\tilde{\mathbf{V}}_{a}-\tilde{\mathbf{V}}_{w}\mathbf{S}_{w}^{-1}\mathbf{S}_{a}\right)$ (F.9)
$=\left(\mathbf{S}_{w}\tilde{\mathbf{V}}_{x}-\tilde{\mathbf{V}}_{w}\mathbf{S}_{x}\right)^{-1}\left(\tilde{\mathbf{V}}_{w}\mathbf{S}_{a}-\mathbf{S}_{w}\tilde{\mathbf{V}}_{a}\right).$ (F.10)
Example 5.
For the regional model by Redding-Rossi-Hansberg-ARE2017, we have
$s_{i}(\bm{x},\bm{w},\bm{a})=w_{i}x_{i}-\sum_{j\in\mathcal{I}}\dfrac{x_{i}a_{i}w_{i}^{1-\sigma}\phi_{ij}}{\sum_{k\in\mathcal{I}}x_{k}a_{k}w_{k}^{1-\sigma}\phi_{kj}}e_{j}=0.$ (F.11)
So, $\mathbf{S}_{a}=-\left(\operatornamewithlimits{diag}[\mathbf{M}\bm{e}]-\mathbf{M}\operatornamewithlimits{diag}[\bm{e}]\mathbf{M}^{\top}\right)\operatornamewithlimits{diag}[\bm{a}]^{-1}=-\frac{\bar{e}}{\bar{a}}\left(\mathbf{I}-\mathbf{D}^{2}\right)$. See Section F.2.2. ∎
Example 6.
For the Krugman model, we have
$s_{i}(\bm{x},\bm{w},\bm{a})=w_{i}x_{i}-\sum_{j\in\mathcal{I}}\dfrac{x_{i}w_{i}^{1-\sigma}\phi_{ij}}{\sum_{k\in\mathcal{I}}x_{k}w_{k}^{1-\sigma}\phi_{kj}}e(w_{j},x_{j},a_{j})=0$ (F.12)
where $e$ maps the tuple $(w_{j},x_{j},a_{j})$ to the regional expenditure. Then, we have $\mathbf{S}_{a}=-\mathbf{M}\mathbf{E}_{a}$, or $\mathbf{S}_{a}=-\epsilon_{a}\mathbf{D}$ at $\bar{\bm{x}}$ where $\epsilon_{a}=\frac{\partial e(\bar{x},\bar{w},\bar{a})}{\partial a_{i}}$. See Section F.2.1. ∎
F.2   Model-specific derivations
We provide omitted derivations of the gain functions $\Omega$, as defined in Section A.2, for the examples in the main text. For derivations for other models mentioned in Sections 2.5 and 2.6, see Akamatsu-et-al-DP2017, an earlier draft of the current paper.
F.2.1   Krugman-JPE1991 model
There are two types of workers, mobile and immobile, and their total masses are $1$ and $L$, respectively. $\bm{x}\equiv(x_{i})_{i\in\mathcal{I}}$ is the distribution of mobile workers. Each worker supplies one unit of labor inelastically.
There are two industrial sectors: agriculture (abbreviated as A) and manufacturing (abbreviated as M). The A-sector is perfectly competitive and a unit input of immobile labor is required to produce one unit of goods. The M-sector follows Dixit–Stiglitz monopolistic competition. M-sector goods are horizontally differentiated and produced under increasing returns to scale using mobile labor as the input. The goods of both sectors are transported. Transportation of A-sector goods is frictionless, while that of M-sector goods is of an iceberg form. For each unit of M-sector goods transported from region $i$ to $j$, only the proportion $1/\tau_{ij}$ arrives, where $\tau_{ij}>1$ for $i\neq j$ and $\tau_{ii}=1$.
All workers have an identical preference for both M- and A-sector goods. The utility of a worker in region $i$ is given by a two-tier form. The upper tier is Cobb–Douglas over the consumption of A-sector goods $C_{i}^{\mathrm{A}}$ and that of M-sector constant-elasticity-of-substitution (CES) aggregate $C_{i}^{\mathrm{M}}$ with $\sigma>1$
$C_{i}^{\mathrm{M}}\equiv\Big{(}\sum_{j\in\mathcal{I}}\int_{0}^{n_{j}}q_{ji}(\xi)^{\frac{\sigma-1}{\sigma}}\mathrm{d}\xi\Big{)}^{\frac{\sigma}{\sigma-1}},$ (F.13)
that is, $u_{i}=(C_{i}^{\mathrm{M}})^{\mu}(C_{i}^{\mathrm{A}})^{1-\mu}$ where $\mu\in(0,1)$ is the constant expenditure of the latter. With free trade in the A-sector, the wage of the immobile worker is equalized, and we normalize it to unity by taking A-sector goods as the numéraire. Consequently, region $i$’s expenditure on the M-sector goods is given by $e_{i}=\mu(w_{i}x_{i}+l_{i})$ where $l_{i}$ denotes the mass of immobile workers in region $i$.
In the M-sector, to produce $q$ units, a firm requires $\alpha+\beta q$ units of mobile labor. Profit maximization of firms yields the price of differentiated goods produced in region $i$ and exported to $j$ as $p_{ij}=\frac{\sigma\beta}{\sigma-1}w_{i}\tau_{ij}$, which in turn determines gravity trade flow from $j$ to $i$. That is, when $X_{ij}$ denotes the price of M-sector goods produced in region $i$ and sold in region $j$, $X_{ij}=m_{ij}e_{j}$ where the share $m_{ij}\in(0,1)$ is defined by Eq. F.6 with $\phi_{ij}\equiv\tau_{ij}^{1-\sigma}$. The proximity matrix is thus $[\phi_{ij}]=[\tau_{ij}^{1-\sigma}]$.
Given $\bm{x}$, we determine the market wage $\bm{w}\equiv(w_{i})_{i\in\mathcal{I}}$ by the M-sector product market-clearing, zero-profit, and mobile labor market-clearing conditions. These conditions are summarized by the trade balance $w_{i}x_{i}=\sum_{j\in\mathcal{I}}X_{ij}$, or Eq. F.5 with $e(x_{i},w_{i})=\mu(w_{i}x_{i}+l_{i})$. By adding up Eq. F.5 for the Krugman model, we see $\sum_{i\in\mathcal{I}}w_{i}x_{i}=\frac{\mu}{1-\mu}L$, which constrains the total income of mobile workers at any configuration $\bm{x}$. The existence and uniqueness of the solution for Eq. F.5 follow from standard arguments (e.g., Facchinei-Pang-Book2007). Given the solution $\bm{w}(\bm{x})$ of Eq. F.5, we have the indirect utility of mobile workers, which is given by $v_{i}=\Delta_{i}^{\frac{\mu}{\sigma-1}}w_{i}$, where $\Delta_{i}\equiv\sum_{k\in\mathcal{I}}x_{k}w_{k}^{1-\sigma}d_{ki}$.
Let $l_{i}=l\equiv\frac{L}{N}$ for all $i\in\mathcal{I}$. We have
$\nabla\log\bm{v}(\bar{\bm{x}})$ $=\dfrac{\mu}{\sigma-1}\mathbf{M}^{\top}\operatornamewithlimits{diag}[\bm{x}]^{-1}-\mu\mathbf{M}^{\top}\operatornamewithlimits{diag}[\bm{w}]^{-1}\mathbf{W}_{x}+\operatornamewithlimits{diag}[\bm{w}]^{-1}\mathbf{W}_{x}$ (F.14)
$=\frac{1}{\bar{x}}\dfrac{\mu}{\sigma-1}\mathbf{D}+\frac{1}{\bar{w}}\left(\mathbf{I}-\mu\mathbf{D}\right)\mathbf{W}_{x},$ (F.15)
where Eq. F.4 and Eq. F.8 give $\mathbf{W}_{x}$. By plugging $\delta=\frac{\mu(\bar{w}\bar{x}+l)}{\bar{w}\bar{x}}=1$ and $\epsilon_{x}=\epsilon_{w}=\mu$ to Eq. F.8,
$\mathbf{W}_{x}=\dfrac{\bar{w}}{\bar{x}}\left(\sigma\mathbf{I}-\mu\mathbf{D}-(\sigma-1)\mathbf{D}^{2}\right)^{-1}\left(\mu\mathbf{D}-\mathbf{D}^{2}\right).$ (F.16)
Then, Eq. F.15 and Eq. F.16 imply
$\mathbf{V}=\bar{x}\nabla\log\bm{v}(\bar{\bm{x}})=\frac{\mu}{\sigma-1}\mathbf{D}+\left(\mathbf{I}-\mu\mathbf{D}\right)\left(\sigma\mathbf{I}-\mu\mathbf{D}-(\sigma-1)\mathbf{D}^{2}\right)^{-1}\left(\mu\mathbf{D}-\mathbf{D}^{2}\right),$ (F.17)
or equivalently, $\mathbf{V}=\Omega(\mathbf{D})$ where
$\Omega(\Theta)$ $=\underbrace{\frac{\mu}{\sigma-1}\Theta}_{\text{(a)}}+\underbrace{(1-\mu\Theta)}_{\text{(b)}}\underbrace{\left(\frac{1}{\sigma}\right)\frac{\mu\Theta-\Theta^{2}}{1-\frac{\mu}{\sigma}\Theta-\frac{\sigma-1}{\sigma}\Theta^{2}}}_{\text{(c)}}.$ (F.18)
From Eq. F.18 we have $\mathbf{V}=\Omega^{\flat}(\mathbf{D})^{-1}\Omega^{\sharp}(\mathbf{D})$, where we define
$\Omega^{\sharp}(\Theta)\equiv\mu\left(\dfrac{1}{\sigma-1}+\dfrac{1}{\sigma}\right)\Theta-\left(\dfrac{\mu^{2}}{\sigma-1}+\dfrac{1}{\sigma}\right)\Theta^{2},$ (F.19)
$\Omega^{\flat}(\Theta)\equiv 1-\frac{\mu}{\sigma}\Theta-\frac{\sigma-1}{\sigma}\Theta^{2}.$ (F.20)
Remark 5.
Using the Krugman model as an example, we discuss how economic forces in a model are embedded in $\Omega$ and $\Omega^{\sharp}$. We recall that positive (negative) terms in $\Omega$ represent agglomeration (dispersion) forces. In Eq. F.18, (a) corresponds to the elasticity of price index with respect to agents’ spatial distribution $\bm{x}$, (b) to the elasticity of payoff with respect to nominal wage $\bm{w}$, (c) to the elasticity of wage with respect to agents’ spatial distribution. Here, (a) and the second term in (b) corresponds to the so-called cost-of-living effect through price index; (a) is positive, i.e., it is an agglomeration force, as the price index in a region becomes lower when more agents (firms) locate geographically close regions; the second term in (b) (i.e., $-\mu\Theta$) is negative because higher wage in a region implies higher goods prices in its nearby regions. Also, (b) as a whole is positive, meaning that the payoff of a region is increasing in wages even with the negative effect through price index. The last component (c) includes both positive and negative terms; in its numerator, the first term ($\mu\Theta$) is demand linkage where firms’ profits rise when they are close to regions with high total income, and the second term ($-\Theta^{2}$) is the market-crowding effect due to competition between firms. The sign of (c) is $\Theta$-dependent; for example, it is negative when $\Theta$ is high ($\phi$ is low) and positive otherwise. The denominator of (c) represents the general equilibrium effects through the so-called short-run equilibrium condition under given $\bm{x}$, i.e., Eq. F.5. As $\Omega^{\sharp}$ is obtained by combining these components and collecting terms according to the order of $\Theta$, these economic forces affect both the first- and second-order coefficients of $\Omega^{\sharp}$. Concretely, in Eq. F.19, $\frac{\mu}{\sigma-1}\Theta$ comes from (a), $\frac{\mu}{\sigma}\Theta$ comes from (b) $\times$ (c), $-\frac{1}{\sigma}\Theta^{2}$ comes from (b) $\times$ (c), and $-\frac{\mu^{2}}{\sigma-1}\Theta^{2}=-\left(\frac{\mu^{2}}{\sigma}+\frac{\mu^{2}}{\sigma(\sigma-1)}\right)\Theta^{2}$ comes from all three components while its leading term $-\frac{\mu^{2}}{\sigma}\Theta^{2}$ comes from (b) $\times$ (c). Thus, by considering $\Omega^{\sharp}$ for a model, one can examine the net effect of all economic forces in the model at once, and the net effect is decomposed according to its spatial scale (i.e., the order of $\Theta$). ∎
Remark 6.
In Fig. C.3, we set $\mu=0.5$, $\sigma=10$, and $L=8$. ∎
Remark 7.
To obtain $\Omega^{\natural}$ for $\bm{l}=(l_{i})_{i\in\mathcal{I}}$, we evaluate $\mathbf{V}_{l}=-\tilde{\mathbf{V}}_{w}\mathbf{S}_{w}^{-1}\mathbf{S}_{l}$ as $\mathbf{A}=\frac{l}{\bar{v}}\mathbf{V}_{l}$. From Example 6, $\mathbf{S}_{l}=-\mu\mathbf{D}$. Also, $\tilde{\mathbf{V}}_{w}=\bar{v}\frac{\partial}{\partial\bm{w}}\log\bm{v}(\bar{\bm{x}})=\frac{\bar{v}}{\bar{w}}(\mathbf{I}-\mu\mathbf{D})$ and $\tilde{\mathbf{V}}_{l}=\bm{0}$. Thus,
$\Omega^{\natural}(\Theta)=c\frac{\Theta(1-\mu\Theta)}{\Omega^{\flat}(\Theta)}>0$ (F.21)
where $c=\frac{l}{\bar{w}}\frac{\mu}{\sigma}=\frac{1-\mu}{\sigma}\bar{x}>0$. It then follows that
$\delta(\Theta)=-\frac{\bar{x}}{\bar{a}}\frac{\Omega^{\natural}(\Theta)}{\Omega(\Theta)}=-\frac{c\bar{x}}{\bar{a}}\frac{\Theta(1-\mu\Theta)}{\Omega^{\sharp}(\Theta)}.$ (F.22)
Straightforward algebra verifies that $\delta^{\prime}(\Theta)<0$ if $\Omega^{\sharp}(\Theta)>0$. ∎
F.2.2   Helpman-Book1998 and Redding-Sturm-AER2008 model
Helpman-Book1998 removed the A-sector in the Krugman model and assumed that all workers are mobile. Instead of the A-sector, the model introduces the housing (abbreviated as H) sector. Each region $i$ is endowed with a fixed stock $a_{i}$ of housing. Workers’ preference is Cobb–Douglas of M-sector CES aggregate $C_{i}^{\mathrm{M}}$ and H-sector goods $C_{i}^{\mathrm{H}}$, $u_{i}=(C_{i}^{\mathrm{M}})^{\mu}(C_{i}^{\mathrm{H}})^{\gamma}$, where $\mu\in(0,1)$ is the expenditure share of the former and $\gamma=1-\mu\in(0,1)$ is that for the latter. There are two variants for assumptions on how housing stocks are owned: public landownership (abbreviated as PL) and local landownership (LL). The original formulation by Helpman-Book1998 supposes housing stocks are equally owned by all workers (i.e., PL). The income of a worker in region $i$ is the sum of the wage and an equal dividend $r>0$ of the total rental revenue in the economy. However, ottaviano2002agglomeration, Murata-Thisse-JUE2005, and Redding-Sturm-AER2008 assumed that housing stocks are locally owned (i.e., LL). The income of a worker in region $i$ is the sum of the wage and an equal dividend of rental revenue in each region. In fact, the model by Redding-Sturm-AER2008 is the LL version of the Helpman model.
Regarding the market equilibrium conditions, the only difference from the Krugman model is regional expenditure $e_{i}$ on M-sector goods in each region:
[PL] $e_{i}=\mu\left(w_{i}+r\right)x_{i},$ (F.23)
[LL] $e_{i}=w_{i}x_{i}.$ (F.24)
Also, the market wage is given as the solution for Eq. F.5. For the PL case, we set $r=1$ for normalization. For the LL case, $\bm{w}(\bm{x})$ is uniquely given up to normalization. The indirect utility function is, with $\Delta_{i}\equiv\sum_{j\in\mathcal{I}}x_{j}w_{j}^{1-\sigma}\phi_{ji}$ and $r>0$,
[PL] $v_{i}=\left(\frac{x_{i}}{a_{i}}\right)^{-\gamma}(w_{i}+r)^{\mu}\Delta_{i}^{\frac{\mu}{\sigma-1}},$ (F.25)
[LL] $v_{i}=\left(\frac{x_{i}}{a_{i}}\right)^{-\gamma}w_{i}^{\mu}\Delta_{i}^{\frac{\mu}{\sigma-1}}.$ (F.26)
Let $a_{i}=1$ for all $i\in\mathcal{I}$. We compute that
$\mathbf{V}=\bar{x}\left(\frac{\mu}{\sigma-1}\mathbf{M}^{\top}\operatornamewithlimits{diag}[\bm{x}]^{-1}+\hat{\mathbf{V}}_{w}\mathbf{W}_{x}-\gamma\operatornamewithlimits{diag}[\bm{x}]^{-1}\right),$ (F.27)
[…]
and $\mathbf{M}$ is defined by Eq. F.6. For the PL case, we obtain $\mathbf{V}=\Omega(\mathbf{D})$ with
$\Omega(\Theta)=-\gamma+\frac{\mu}{\sigma-1}\Theta+\frac{\mu\left(\mu-\Theta\right)\Theta(1-\Theta)}{\sigma-\mu\Theta-(\sigma-1)\Theta^{2}}$ (F.30)
where we compute $\mathbf{W}_{x}$ from Eq. F.7 with $\zeta=\frac{\mu(\bar{w}+1)}{\bar{w}}$, $\epsilon_{x}=\frac{\bar{w}+1}{\bar{w}}$, $\epsilon_{w}=\mu$; we note that $\frac{\bar{w}}{\bar{w}+1}=\mu$ under our normalization. Thus, for the PL case, we can choose $\Omega^{\sharp}$ and $\Omega^{\flat}$ that satisfy $\mathbf{V}=\Omega^{\flat}(\mathbf{D})^{-1}\Omega^{\sharp}(\mathbf{D})$ as follows:
$\Omega^{\sharp}(\Theta)\equiv-\gamma+\left(\frac{\mu}{\sigma-1}+\dfrac{\mu(\mu+\gamma)}{\sigma}\right)\Theta-\left(\frac{\mu^{2}}{\sigma-1}+\dfrac{\mu+\gamma}{\sigma}-\gamma\right)\Theta^{2},$ (F.31)
$\Omega^{\flat}(\Theta)\equiv 1-\dfrac{\mu}{\sigma}\Theta-\dfrac{\sigma-1}{\sigma}\Theta^{2}$ (F.32)
where we recall $\mu+\gamma=1$.
Remark 8.
As in the Krugman model (Remark 5), one can identify the correspondence between the coefficients in $\Omega$ Eq. F.31 and those in $\Omega^{\sharp}$ Eq. F.30. For example, local dispersion force $-\gamma$ in Eq. F.30 affects not only constant term in $\Omega^{\sharp}(\cdot)$ but also the first- and the second-order terms of $\Theta$. Other forces can be understood similarly to the Krugman model (Remark 5). ∎
For the LL case, $\mathbf{W}_{x}$ is given in Example 4 and we obtain
$\Omega(\Theta)=-\gamma+\frac{\mu}{\sigma-1}\Theta+\frac{\mu(1-\Theta)\Theta}{\sigma+(\sigma-1)\Theta}.$ (F.33)
We can choose $\Omega^{\sharp}$ and $\Omega^{\flat}$ as follows:
$\Omega^{\sharp}(\Theta)\equiv-\gamma+\left(\frac{\mu}{\sigma-1}+\frac{\mu+\gamma}{\sigma}-\gamma\right)\Theta,$ (F.34)
$\Omega^{\flat}(\Theta)\equiv 1+\dfrac{\sigma-1}{\sigma}\Theta.$ (F.35)
Remark 9.
Definition 2 classifies canonical models based on the spatial scale of the “effective” dispersion forces. The Helpman model with public landownership Eq. F.31 has global dispersion forces because $c_{2}<0$ under Assumption E. However, unlike the Krugman model, the global dispersion forces in the Helpman model are not “effective” in the sense that, under any admissible values of $\mu$ and $\sigma$, this force does not stabilize the uniform distribution for any level of transport costs. If we drop the local dispersion force $c_{0}<0$ from $\Omega^{\sharp}$, we have $c_{1}\Theta+c_{2}\Theta^{2}>0$ for all $\Theta$ and $\bar{\bm{x}}$ is always unstable. Thus, the only dispersion force in the Helpman model that can stabilize $\bar{\bm{x}}$ is its local dispersion force. ∎
Remark 10.
Equilibrium is unique when $\gamma\sigma=(1-\mu)\sigma>1$ (Redding-Sturm-AER2008). For both PL and LL, it implies that $\Omega^{\sharp}(\Theta)<0$ for all $\Theta\in(0,1)$. ∎
Remark 11.
The regional model in §3 of Redding-Rossi-Hansberg-ARE2017 is a variant of the Helpman model with LL, in which variable input of mobile labor depends on region $i$ (i.e., productivity differs across regions). The cost function of firms in region $i$ is $C_{i}(q)=w_{i}(\alpha+\beta_{i}q)$. The market equilibrium condition for this case is, with $a_{i}\equiv\beta_{i}^{1-\sigma}>0$, given by
$s_{i}(\bm{x},\bm{w})=w_{i}x_{i}-\sum_{j\in\mathcal{I}}\dfrac{x_{i}a_{i}w_{i}^{1-\sigma}\phi_{ij}}{\sum_{k\in\mathcal{I}}x_{k}a_{k}w_{k}^{1-\sigma}\phi_{kj}}w_{j}x_{j}=0.$ (F.36)
The payoff function is given by Eq. F.29 with $\Delta_{i}=\sum_{k\in\mathcal{I}}x_{k}a_{k}w_{k}^{1-\sigma}\phi_{ki}$.
From Example 5, $\mathbf{S}_{a}=-\frac{\bar{w}\bar{x}}{\bar{a}}\left(\mathbf{I}-\mathbf{D}\right)\left(\mathbf{I}+\mathbf{D}\right)$ as $\bar{e}=\bar{w}\bar{x}$. Also, we have $\tilde{\mathbf{V}}_{w}=\frac{\bar{v}}{\bar{w}}\mu(\mathbf{I}-\mathbf{D})$, $\tilde{\mathbf{V}}_{a}=\frac{\bar{v}}{\bar{a}}\frac{\mu}{\sigma-1}\mathbf{D}$, and $\mathbf{S}_{w}=\sigma\bar{x}\Omega^{\flat}(\mathbf{D})$. As $\mathbf{V}_{a}=\tilde{\mathbf{V}}_{a}-\tilde{\mathbf{V}}_{w}\mathbf{S}_{w}^{-1}\mathbf{S}_{a}$ and $\mathbf{A}=\frac{\bar{a}}{\bar{v}}\mathbf{V}_{a}=\Omega^{\natural}(\mathbf{D})$, we compute
$\Omega^{\natural}(\Theta)=c\frac{(\sigma-1)+\sigma\Theta}{\Omega^{\flat}(\Theta)}>0$ (F.37)
where $c\equiv\frac{\bar{v}}{\bar{a}}\frac{\mu}{\sigma}>0$. This in turn implies
$\delta(\Theta)=-\frac{\bar{x}}{\bar{a}}\frac{\Omega^{\natural}(\Theta)}{\Omega(\Theta)}=-\frac{c\bar{x}}{\bar{a}}\frac{(\sigma-1)+\sigma\Theta}{\Omega^{\sharp}(\Theta)}$ (F.38)
where $\Omega^{\sharp}(\Theta)$ is that for the LL case Eq. F.35. We have $\delta^{\prime}(\Theta)>0$ for all $\Theta$ whenever $(1-\mu)\sigma>1$. ∎
F.2.3   Pfluger-Suedekum-JUE2008 model
The Pflüger–Südekum model builds on Pfluger-RSUE2003 and introduces the housing sector (again denoted by H). In this model, a quasi-linear upper-tier utility is assumed: $u_{i}=C_{i}^{\mathrm{A}}+\mu\log C_{i}^{\mathrm{M}}+\gamma\log C_{i}^{\mathrm{H}}$. The production cost for a firm in $i\in\mathcal{I}$ is $\alpha w_{i}+\beta q$. Then, $\bm{w}$ is given as follows:
$w_{i}=\dfrac{\mu}{\sigma}\sum_{j\in\mathcal{I}}\dfrac{\phi_{ij}}{\sum_{k\in\mathcal{I}}\phi_{kj}x_{k}}(x_{j}+l_{j}).$ (F.39)
The indirect utility of a mobile worker in region $i$ is
$v_{i}(\bm{x})=\frac{\mu}{\sigma-1}\ln[\Delta_{i}]-\gamma\ln\frac{x_{i}+l_{i}}{a_{i}}+w_{i},$ (F.40)
where $\Delta_{i}=\sum_{j\in\mathcal{I}}\phi_{ji}x_{j}$, and $l_{i}$ and $a_{i}$ denote the mass of immobile workers and amount of housing stock in region $i$, respectively. The nominal wage in region $i$ is given by Eq. F.39. Let $l_{i}=l$ and $a_{i}=a$ for all $i$. Then, we see that $\mathbf{V}=\frac{1}{\bar{v}}\Omega^{\sharp}(\mathbf{D})$ with
$\Omega^{\sharp}(\Theta)$ $=-\dfrac{\gamma}{1+L}+\mu\left(\frac{1}{\sigma-1}+\frac{1}{\sigma}\right)\Theta-\frac{\mu}{\sigma}(1+L)\Theta^{2}.$ (F.41)
Remark 12.
In Figs. 2A and 2B, $\mu=0.4$, $\sigma=2.5$, $L=4$, $\gamma=0.5$, and $a_{i}=1$. ∎
F.2.4   Allen-Arkolakis-QJE2014 model
The AA model is a perfectly competitive Armington-IMF1969-based framework with positive and negative local externalities. We consider a discrete-space version of the model and abstract away all exogenous differences across regions. In the model, productivity of a location is proportional to $x_{i}^{\alpha}$ with $\alpha>0$, representing positive externalities. The market equilibrium condition is
$s_{i}(\bm{x},\bm{w})=w_{i}x_{i}-\sum_{j\in\mathcal{I}}\dfrac{w_{i}^{1-\sigma}x_{i}^{\alpha(\sigma-1)}\phi_{ij}}{\sum_{k\in\mathcal{I}}w_{k}^{1-\sigma}x_{k}^{\alpha(\sigma-1)}\phi_{kj}}w_{j}x_{j}=0.$ (F.42)
With market wage $\bm{w}$, we have $v_{i}(\bm{x})=x_{i}^{\beta}w_{i}\Delta_{i}^{\frac{1}{\sigma-1}}$ with $\Delta_{i}\equiv\sum_{k\in\mathcal{I}}w_{k}^{1-\sigma}x_{k}^{\alpha(\sigma-1)}\phi_{ki}$; $x_{i}^{\beta}$ ($\beta<0$) represents negative externalities from congestion. We have $\mathbf{V}=\Omega^{\flat}(\mathbf{D})^{-1}\Omega^{\sharp}(\mathbf{D})$ with
$\Omega^{\sharp}(\Theta)=\left(\alpha+\beta-\frac{1+\alpha}{\sigma}\right)+\left(\alpha+\beta+\frac{1-\beta}{\sigma}\right)\Theta,$ (F.43)
$\Omega^{\flat}(\Theta)=\left(\sigma+(\sigma-1)\Theta\right)\left(1-\Theta\right).$ (F.44)
It is also instructive to consider the case with no externalities ($\alpha=0$ and $\beta=0$). Then, the model reduces to the Armington-IMF1969 trade model, and we have $\omega=-\frac{1}{\sigma(\sigma+(\sigma-1)\Theta)}<0$. Thus, the intrinsic working of general equilibrium effects induced by love for variety in the Armington model is in creating a dispersion force, and it is global in the sense that the force depends on $\Theta$. However, in the context of net agglomeration incentive at the symmetry with nonzero spillovers, this force is mainly related to the denominator $\Omega^{\flat}$, and $\Omega^{\sharp}$ succinctly summarizes the net trade-off between agglomeration and dispersion forces that govern the stability of $\bar{\bm{x}}$.
Figure F.1 provides the classification of spatial patterns and their stability under Assumption C. Figure F.1 can be seen as a refinement of Figure I in Allen-Arkolakis-QJE2014 under Assumption C. A sufficient condition for the equilibrium uniqueness is $\alpha+\beta\leq 0$ (Range 3), which means there is net local dispersion force and $\Omega^{\sharp}(\Theta)<0$ for all $\Theta$. In Range 3, $\bar{\bm{x}}$ is the only equilibrium.
Remark 13.
In Fig. C.3, we set $(\alpha,\beta,\sigma)=(0.5,-0.3,6)$. ∎
Figure F.1: Uniqueness and stability of equilibria in the Allen–Arkolakis model.
F.2.5   Krugman-Venables-QJE1995 model
Krugman-Venables-QJE1995 considers intermediate inputs. When we interpret the regional share of manufactured good production as the state variable, the model is Type G. With input share of manufactured good in production being $\alpha$, Fujita-Krugman-Venables-Book1999, Appendix 14.1, derives a concise formula for $\Omega$. Namely, in the case that “agricultural” good is produced by a constant returns technology,
$\Omega(\mathbf{D})=\Omega^{\flat}(\mathbf{D})^{-1}\left(\alpha\left(1+\frac{\sigma-1}{\sigma}\right)\mathbf{D}-(1+\alpha)^{2}\mathbf{D}^{2}\right)$ (F.45)
with an appropriately defined $\Omega^{\flat}$, showing that the model is Type G. Also, if the “agricultural” sector exhibits decreasing return, an extra local dispersion force (a term $c_{0}\mathbf{I}$ such that $c_{0}<0$) emerges, so that the model becomes Type LG (Fujita-Krugman-Venables-Book1999, Section 14.4).
F.2.6   Kucheryavyy-etal-JIE2024
Kucheryavyy-etal-JIE2024 considers a unified framework that nests Allen-Arkolakis-QJE2014 and Krugman-JPE1991 as special cases. For this model in the symmetric two-region economy, the sign of the utility gain $\omega$ coincides with the following function of $\Theta$:
$\Omega^{\sharp}(\Theta)\equiv-(1-\alpha)+(1-(1-\beta)\zeta+\mu)\Theta-(\alpha+(1-(1-\beta)\zeta)\mu)\Theta^{2},$ (F.46)
where the notations follow the original paper except for the transport cost index $\Theta$.
For deriving the above $\Omega^{\sharp}$, we note that, when evaluated at $x=1$, $V^{\prime}(x)$ in Appendix D of Kucheryavyy-etal-JIE2024 is essentially a negative constant multiple of the (net) utility gain, as the stability condition for the uniform distribution in their notation is $V^{\prime}(1)>0$ (see their Appendix B.4).
As seen, the model is Type LG in its most general form. Importantly, the case $(\alpha,\zeta,\mu)=(1,1,\beta)$ corresponds to a Krugman-type model, as we observe Eq. F.46 reduces to a function of the form $c_{1}\Theta-c_{2}\Theta^{2}$. Also, $\alpha<1$ and $\zeta=1$ corresponds to the Allen–Arkolakis framework, for which we confirm $-(1-\alpha)<0$ represents a local dispersion force. It is noted that $\Omega^{\sharp}(1)=(1-\mu)(1-\beta)$. If $\mu<1$, then $\Omega^{\sharp}(1)>0$, so that $\Omega^{\sharp}$ must have one and one zero for $\Theta\in(0,1)$, so that $\bar{\bm{x}}$ is stable only for low transport costs.
References for Supplementary Materials